### v2

# Structured Packing in LLM Training Improves Long Context Utilization

Abstract
Recent advances in long-context Large Language Models (LCLMs) have generated significant interest, especially in applications such as querying scientific research papers. However, their potential is often limited by inadequate context utilization. We identify the absence of long-range semantic dependencies in typical training data as a primary hindrance. To address this, we delve into the benefits of frequently incorporating related documents into training inputs. Using the inherent directory structure of code data as a source of training examples, we demonstrate improvements in perplexity, even for tasks unrelated to coding. Building on these findings, but with a broader focus, we introduce Structured Packing for Long Context (SPLiCe). SPLiCe is an innovative method for creating training examples by using a retrieval method to collate the most mutually relevant documents into a single training context. Our results indicate that SPLiCe enhances model performance and can be used to train large models to utilize long contexts better. We validate our results by training a large $3$B model, showing both perplexity improvements and better long-context performance on downstream tasks.
1 Introduction
Large language models (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023; Bai et al., 2023) have completely transformed the field of AI and natural language processing. Long-context language models (LCLMs) are becoming increasingly popular and have the potential to unlock new capabilities (Anthropic, 2023). However, recent studies (Liu et al., 2023; Tworkowski et al., 2023) have highlighted numerous issues with their context utilization. Even if they have the technical ability to process long context, some models fail in practice, as evidenced by their poor performance even on synthetic retrieval tasks (Li et al., 2023a). Moreover, LLMs are often distracted from retrieving relevant information when provided with multiple documents in the prompt (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023), and typically struggle to utilize information from the middle of their input (Liu et al., 2023). There might be multiple reasons behind these issues, requiring structured research to address them.
In this work, we take a step towards better context utilization in LCLMs. We focus on training data, keeping other components (e.g., the architecture and training objectives) unchanged. The fundamental question then is how to construct training examples such that using long context is beneficial for the model (e.g., it allows for obtaining lower perplexity or acquiring new downstream capabilities) while being non-trivial (like simple copying). We show that including multiple related documents in a training context can be beneficial - not only for evaluating on multiple documents, but also when considering language modeling perplexity on long, naturally occurring, single documents, which our evaluation is aimed for. First, we showcase this using the natural directory structure of the code data. Following that, we propose SPLiCe, Structured Packing for Long Context, that constructs training examples consisting of multiple similar documents selected using a retrieval method of choice (e.g., BM25).
Figure 1: Training examples generated by the standard sampling procedure and proposed SPLiCe method.
We empirically validate SPLiCe on a large-scale code generation dataset (Li et al., 2023b), demonstrating improved perplexity evaluated at larger contexts, as well as a higher performance on TREC (Li & Roth, 2002; Hovy et al., 2001), Qasper, and synthetic task from (Liu et al., 2023). For that, we use the dataset generated by our method to continue pretraining an OpenLLaMA 3B LLM using the methodology described in (Tworkowski et al., 2023), resulting in a long-context language model with improved capabilities. Our contributions can be summarized as follows:
  • We show that structuring training data is a viable way of improving the context utilization of LLMs. We validate this using the natural directory structure of code data.
  • We introduce SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes retrieval (implemented using BM25 or Contriever (Izacard et al., 2022)) to find relevant documents without relying on directory structure. We show that using SPLiCe during training on code matches the model’s performance of the directory-structure method while being more generally applicable.
  • We present a thorough analysis of the design choices of SPLiCe, the number of retrieved documents, and the order in which documents are merged into a training example.
  • We validate SPLiCe by training a large-scale model, OpenLLaMA 3B v2 (Geng & Liu, 2023). We show that SPLiCe leads to improved perplexity on long-context evaluation, improved in-context learning ability and better retrieval performance.
2 Method
SPLiCe is a method of constructing the training examples that allows improvement of large-scale models’ long-context utilization and, consequently, their performance on a range of tasks. Through experimental study, we positively validate that such an approach leads to better long-context capabilities.
Rationale and intuitions
Capturing long-range dependencies is believed to enhance language modeling, e.g. as demonstrated in the study by (Borgeaud et al., 2022) which highlights the advantages of incorporating global dependencies in retrieval-augmented language models. However, in typical datasets, the occurrence of such dependencies is rare (de Vries, 2023) and diminishes as the distance between elements increases, leading to challenges in model learning. Our approach addresses this challenge by focusing on increasing dependency density, achieved by creating training examples that consist of multiple related documents. The rationale of this strategy is reinforced by the work of Levine et al. (2022), which shows that the trained model establishes stronger dependencies between text segments present in the same training example.
In our work, we study the following methods:
Baseline
The standard approach, commonly used in LLM training pipelines (Brown et al., 2020), is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context (also known as example packing).
Repository-level code data
Certain datasets encompass meta-information regarding the data structure, which can be utilized to devise training examples. Take, for instance, a dataset consisting of GitHub repositories (e.g., (Li et al., 2023b)): rather than grouping documents randomly into a single context, we might group documents based on their subdirectory in a repository, thereby crafting a training example with potentially more cross-document dependencies compared to those generated randomly. This methodology, elaborated in (Wu et al., 2022), led to notable perplexity enhancements with augmented context length, suggesting that data orchestrated in such a manner could exhibit valuable long-context training signals. In our study, we strive to quantify the advantages of this methodology against the random packing baseline outlined above while maintaining the set of training documents to eliminate potential confounders.
Figure 2: Visualization of SPLiCe training example generation.
Structured Packing for Long Context (SPLiCe)
The additional structure provided by meta-information is unavailable in most datasets. We propose to overcome this challenge by using a retrieval method to identify related documents. We test two choices of retrieval methods BM25 (Robertson & Zaragoza, 2009; Bassani, 2023) and Contriever-MSMARCO (Izacard et al., 2022)). With those tools at our disposal, we propose SPLiCe, a method that creates training examples by building a tree of documents. It starts from a randomly sampled document and continues in a breadth-first manner, expanding each node with top-$k$ similar documents from the corpus. The final sequence consists of a flattened tree according to a certain traversal strategy (e.g., level-order traversal). An intuitive visualization of SPLiCe is presented in Figures 1 and 2, while a detailed pseudocode is given in Algorithm 1. We note that the general formulation of SPLiCe is flexible. For example, the parameter $k$ allows for interpolation between two modes of retrieval. For $k=1$ SPLiCe creates a path of related examples simulating a long document. For larger values of $k$ our method generates examples that are similar to the ones used by retrieval augmented models. Note that the dataset must be reasonably curated for SPLiCe to work efficiently. For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
Algorithm 1 SPLiCe training example construction
Input:
   $D$: the document corpus,
   $k$: the breadth hyper-parameter
   $L$: Maximum length of the returned training example
Output: training example consisting of concatenated documents
$d_{r}\sim D$triangleright Sample the root document
$D=D\setminus\{d_{r}\}$
$C=d_{r}$
$Q=\texttt{empty queue }$
$Q.\texttt{PUSH}(d_{r})$
while $Q\neq\emptyset$ and $\texttt{len}(C)\leq L$ do
      $d=Q.\texttt{POP}()$
      $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$triangleright Retrieve top $k$ most similar documents to $d$ using a                                                         selected method, e.g., BM25
     for each $d_{i}$ in $d_{1},\ldots,d_{k}$ do
         if $d_{i}\in D$ then $\triangleright$ RETRIEVE uses a precomputed index, and may return                                                     documents that are already in $C$
                  $C=C.\texttt{CONCAT}(d_{i})$triangleright Append $d_{i}$ to $C$
                  $Q.\texttt{PUSH}(d_{i})$
                  $D=D\setminus\{d_{i}\}$
return $\texttt{TRIM}(C,L)$
3 Experiments with medium-scale models
In this section, we conduct a comprehensive examination of the impact of document packing into training examples on long-context performance. We demonstrate that leveraging the inherent repository structure of code data can yield notable perplexity enhancements. Additionally, we establish that our proposed method, SPLiCe, either matches or surpasses this approach, exhibiting broader applicability to non-code data. We stress that we do not aim to surpass the approach to concatenate repository files together, instead we aim to show that SPLiCe is more general and can be used for non-code data. We carry out an ablation study on SPLiCe to scrutinize its design elements and hyperparameters in Appendix C.1. For this segment, due to computational constraints, we engage with $270$M models to ensure a meticulous investigation. In Section 4, we affirm that these findings extend to large-scale models.
3.1 Experimental setup
To investigate the impact of input data organization on enabling language models to leverage extended context, we employ $270$M models following a specific protocol. Initially, we pre-train a $270$M model on $6.3$B tokens from the original RedPajama mixture (TogetherComputer, 2023), using a context size of $2$K. Subsequently, we continue pretraining this model on $1$B tokens with an extended context length of $32$K, employing the Focused Transformer (FoT) objective Tworkowski et al. (2023), on a balanced mixture of the original RedPajama data (TogetherComputer, 2023) and long-context data assembled using the specified method. This approach is driven by practical factors: training with a shorter context expedites the process, and as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023), the model training can be continued on longer documents from the original data source while maintaining the desired long-context perplexity scaling. During the long-context pretraining phase, only the latter half of the mixture is altered, retaining the original RedPajama portion intact. This approach is loosely inspired by (Ouyang et al., 2022) and (Rozière et al., 2023) and aims to prevent the model from overfitting to artificially created long documents. Only the effect of data organization into training examples is tested, keeping other factors like architecture intact. For a detailed description of the model architecture, refer to Appendix A, and for information regarding the training data, see Appendix F.
In the evaluation, we employ a context length of $32$K and measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets. The selection of these evaluation datasets is motivated by their inclusion of documents that stand to benefit from long-context information. Specifically, functions are often re-utilized, and similar terms are employed across papers. The advantage of long context in these datasets has been demonstrated for both arXiv (Chen et al., 2023) and StarCoder (Li et al., 2023b). We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. For further details regarding the chosen datasets and the evaluation setup, refer to Appendix F.
3.2 SPLiCe on code data
In this section, we compare SPLiCe against the Baseline and repository-level code data (Repo) packing methods. We test two retrieval approaches for SPLiCe, the first one is BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), the second one is and Contriever-MSMARCO (Izacard et al., 2022)) and denote them as SPLiCe BM25 and SPLiCe Cont respectively. In SPLiCe BM25 retrieval takes into account the whole content of the document, whereas in SPLiCe Cont we use the first $512$ tokens and Faiss (Johnson et al., 2017b) for fast approximate inner-product search (for details see Appendix H). To make this comparison possible we use code data (required by Repo). We observe, see Table 1, that SPLiCe significantly outperforms the Baseline, and is either slightly better or on par with Repo. This is a promising result, as SPLiCe is more general and can be applied to non-code data (see Section 3.3).
Table 1: We train a $270$M parameter model on a 50/50 mixture of the RedPajama data (organized in a standard way) and code data (organized in one of four ways: SPLiCe BM25, SPLiCe Cont, Repo, Baseline – standard). We evaluate the perplexity of the models on arXiv and various subsets of StarCoder. The boldface denotes the best result within the training mixture. Note that SPLiCe is generally on par or better than Repo and beats the Baseline by a large margin. For detailed results, see Appendix B.
Training Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
Repo 5.47 3.22 2.83 2.24 2.96 3.12
C# SPLiCe BM25 5.52 3.33 2.90 2.46 3.11 3.26
SPLiCe Cont 5.53 3.35 2.91 2.48 3.12 3.27
Baseline 5.65 3.58 3.07 2.65 3.31 3.46
Repo 5.53 3.35 2.91 2.49 3.12 3.27
Python SPLiCe BM25 5.47 3.25 2.53 2.41 3.02 3.17
SPLiCe Cont 5.49 3.28 2.53 2.43 3.03 3.19
Baseline 5.57 3.46 2.62 2.56 3.17 3.32
Repo 5.48 3.27 2.54 2.44 3.03 3.18
To deepen the analysis, we prepared three subsets of the C code and ran the tested methods on each of them. In Table 2, we report the mean and standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results. For details about how the training sets were prepared please refer to Appendix F.
Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. For the C code we use three different subsets and report the mean and standard deviation.
Training Method arXiv Code Code &
Data Python All arXiv
C SPLiCe BM25 5.463 $\pm$ 0.002 2.810 $\pm$ 0.002 2.942 $\pm$ 0.005 3.100 $\pm$ 0.004
SPLiCe Cont 5.477 $\pm$ 0.005 2.824 $\pm$0.001 2.957 $\pm$ 0.006 3.115 $\pm$ 0.006
Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
Repo 5.474 $\pm$ 0.007 2.827 $\pm$ 0.006 2.958 $\pm$ 0.009 3.115 $\pm$ 0.009
3.3 SPLiCe on webtext data
The major advantage of our method over Repo is that SPLiCe can be applied to any textual data. Table 3 shows that using our method the model outperforms the baseline when trained on a $/50$ 50 mixture of RedPajama (packed in the standard way) and Wikipedia/StackExchange (prepared using SPLiCe). Notably, we observe that the model trained on StackExchange data significantly outperforms the one trained on C in the arXiv evaluation. Despite the non-code nature of StackExchange we still note perplexity improvements on code. Unsurprisingly, the perplexity on code is not as good as when training directly on code data. We conjecture that better data mixtures could bring further benefits.
Table 3: Perplexity results for training on a $/50$ 50 data mixture of RedPajama and one of {StackExchange,C} organized using mentioned methods. Note that when evaluating on arXiv the model trained with SPLiCe organized StackExchange outperforms the one trained on C.
Training Method arXiv Code Code &
data Haskell Python CUDA All arXiv
Wikipedia SPLiCe BM25 5.64 3.82 3.26 2.87 3.55 3.68
SPLiCe Cont 5.65 3.87 3.30 2.92 3.59 3.72
Baseline 5.73 3.97 3.37 3.00 3.68 3.81
StackExchange SPLiCe BM25 5.07 3.88 3.32 2.89 3.60 3.69
SPLiCe Cont 5.09 3.91 3.35 2.93 3.63 3.73
Baseline 5.14 3.94 3.36 2.94 3.65 3.74
C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
3.4 SPLiCe with different context extension approaches
Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension. FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. This allows for computationally efficient long-context fine-tuning. To show that our method works more generally, we have also evaluated it using more standard approaches. That is, we checked the effectiveness of the method when context extension is done in all layers, and parameters of Rotary Positional Encodings are either adjusted as in CodeLlama (Rozière et al., 2023), adjusted with YaRN (Peng et al., 2023), or left without changes. Table 4 shows the results.
Table 4: Perplexity results for training on a $/50$ 50 data mixture of RedPajama, and C# organized using the mentioned methods. We check how much SPLiCe can help when fine-tuning for the larger context using methods different than FoT (Tworkowski et al., 2023). That is, we test the effectiveness of SPLiCe when the context is extended to $16K$ in all layers and parameters of positional embedding are either adjusted as in CodeLlama (Rozière et al., 2023), with YaRN (Peng et al., 2023), or left without changes. The results indicate that SPLiCe still brings perplexity improvements in all of these setups. For YaRN we have used the parameters from (Peng et al., 2023) and adjusted the scale factor appropriately.
RoPe scale Training Method arXiv Code Code &
method data Haskell Python CUDA All arXiv
Naive C# SPLiCe BM25 6.25 4.84 3.55 2.84 3.72 3.88
Baseline 6.33 5.03 3.66 2.96 3.85 4.00
Repo 6.25 4.87 3.56 2.85 3.74 3.89
CodeLlama C# SPLiCe BM25 5.74 4.28 3.22 2.53 3.34 3.49
Baseline 5.76 4.37 3.27 2.58 3.40 3.55
Repo 5.74 4.28 3.22 2.54 3.35 3.50
YaRN C# SPLiCe BM25 5.77 4.32 3.24 2.55 3.37 3.52
Baseline 5.79 4.42 3.29 2.61 3.44 3.58
Repo 5.77 4.32 3.24 2.56 3.38 3.53
4 Large-scale models
In this section, we show that SPLiCe can improve the long context performance of large-scale language models. To this end, we use $3$B parameter models and tasks that test in-context learning, question answering, and information retrieval capabilities. We provide perplexity results in Appendix E
4.1 Experimental setup
We continue training OpenLLaMA 3B v2 (Geng & Liu, 2023; Geng, 2023) on the dataset defined in Section 3.1 - that is, a 50/50 mixture of RedPajama prepared in a standard way and C prepared using a method of choice. Specifically, we train two models with $32K$ context length for the same number of tokens, 5.4B, which is $21$k training steps, on two datasets: one generated using SPLiCe, and the other one using the Baseline, see Section 2 for definitions. We use a batch size of 256K tokens per step, and learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023). We evaluate the perplexity on a mixture of long documents sourced from CommonCrawl, arXiv, and StarCoder (Li et al., 2023b). Otherwise, we follow the protocol described in Section 3.1.
4.1.1 In context learning
To evaluate the improvements of the model’s in-context learning ability we use TREC (Li & Roth, 2002; Hovy et al., 2001) question classification task. We provide the SPLiCe trained model with $190$ to $1560$ in-context examples and compare the prediction accuracy against the Baseline model. We sample the few-shot examples using 50 different random seeds and show the histogram of accuracy improvements in Figure 3 and additional results in Table 4 and Figure 7.
Figure 3: Histogram of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of $1560$ elements (roughly $32$k tokens) randomly sampled (without replacement) from the training subset of TREC.
4.2 Standardized Comparison Over Long Language Sequences
We show that SPLiCe can improve model performance on Qasper (Dasigi et al., 2021; Shaham et al., 2022). Qasper tests the model’s ability to answer questions regarding research papers provided as a part of the input. The evaluated models were trained on a mixture of RedPajama prepared in a standard way and C prepared either in a standard way (Baseline) or using SPLiCe. Note that despite SPLiCe using only code data here, we still see improvements in score on non-code, see Table 5.
Table 5: Zero-shot performance on the validation subset of Qasper (Dasigi et al., 2021) and few-shot performance on TREC. For Qasper we use the implementation from Language Model Evaluation Harness (Gao et al., 2021). In the Harness implementation, yes/no questions are evaluated separately from open questions. Note that despite using SPLiCe for code data only we still have improvements in non-code tasks. For TREC we average results across $50$ sets of in-context examples. Note that for TREC SPLiCe model is better for almost all choices of in-context examples (see Figures 3 and 7).
Task Context length 3B Baseline 3B SPLiCe
TREC 32K 73.9 $\pm$3.9 79.3 $\pm$ 2.9
16K 68.9 $\pm$5.9 76.9 $\pm$ 3.1
8K 66.5 $\pm$6.2 75.8 $\pm$ 3.5
4K 65.1 $\pm$6.1 72.8 $\pm$ 4.9
Qasper F1 32K 23.1 23.9
4K 18.6 18.7
4.3 Key Retrieval Performance
Recent works indicated numerous issues related to long-context utilization. In particular, (Liu et al., 2023, Figure 7) proposed a benchmark retrieval task to measure the ability of language models to retrieve information from a long context. Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key. Even though very simple, the task proved to be challenging for many open-source language models.
In Figure 4 we present the key retrieval performance of the models trained with SPLiCe and Baseline. The dictionary contains $75$ pairs, which is about 6k tokens. The performance depends heavily on the key position (which is also the case for other models, as indicated in (Liu et al., 2023)). The keys near the beginning are much harder to retrieve, and SPLiCe improves the accuracy more than two times.
Figure 4: Key-value retrieval performance on a dictionary of $75$ key-value pairs ($\approx$ 6k tokens). The model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions. Both models achieve $100$% accuracy near the end because the desired key is in the local context (below $2048$ tokens). The details about this task can be found in Appendix D.
5 Related work
Some recent studies have focused on enhancing training examples to improve the performance of language models. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include ”intrinsic tasks”. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022).
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of this approach in large-scale language modeling, specifically with OpenLLaMA 3B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code.
6 Limitations and future work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method can viewed as a general framework for organizing the documents into training examples (i.e. to achieve the repository-level grouping, it suffices to replace the $\texttt{RETRIEVE}(d,k)$ method with one that returns files from the closest catalog in the repository). This opens multiple further research avenues.
In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Contriever-MSMARCO). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, how the ratio of related and unrelated documents) affects performance for future work. Note that as BM25 is a text-statistic based retriever, it is not guaranteed that the documents returned by it are semantically related. Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
Another, avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
Our method requires further studies to understand its scaling properties beyond the range presented in the paper. In particular, we leave scaling the context length to hundreds of thousands of tokens and experiments with models having more than $3$B parameters for future work.
Our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). This methodology is similar to the approach used in (Wu et al., 2022) and allows extension of the model context with relatively small (compared to other methods) computational cost. Briefly speaking, it boils down to choosing a subset of attention layers for the context extension. We have additionally tested three popular context extension methods on a medium-scale (Naive, YaRN, and CodeLlama). We leave studies of SPLiCe with other context extension methods as future work.
One of the approaches to training long-context language models is to use long conversational data (Li et al., 2023a). This approach is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, and answers, etc.) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
In order to prepare the training data SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
Using highly correlated samples has the potential to result in training instability. However, during our experiments with StarCoder, Wikipedia, and StackExchange data we have noted no performance degradation. We leave the study of how SPLiCe integrates with different types of data for the future. In particular, in our studies used datasets were reasonably deduplicated.
7 Conclusions
In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves the perplexity of on language modeling task and improves the performance on downstream tasks. We further show that SPLiCe can be used to train large-scale models, resulting in a model with improved long-context utilization. Besides its direct usefulness, we believe that our work indicates multiple interesting research directions on how to improve the performance of long-context language models with structured data.
8 Reproducibility
To ensure the reproducibility of our results we provide the details about how the data was prepared in Appendix F, the pseudocode of our method in Algorithm 1 and details about the trained models in Appendix A. We also attach the source code in the supplementary material.
Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al., 2023), and the architectural details can be found in Table 6. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al., 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al., 2021). For context extension, we use Focused Transformer (Tworkowski et al., 2023). Table 7 presents the details about both standard and long-context pretraining.
Table 6: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
Parameter/Model Size 270M 3B
Vocab Size 32000 32000
Embedding Size 1024 3200
Num Attention Layers 12 26
Num Attention Heads 8 32
Head Size 128 100
MLP Hidden Size 4096 8640
FoT Context Extension Layers [6, 9] [6, 12, 18]
Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B parameter OpenLLaMAv2 model (Geng, 2023).
Stage Parameter/Model Size 270M 3B
Pretraining Context 2K 2K
Tokens 6.3B 1T
Long Context Pretraining Context 32K 32K
Batch Size 128 128
Start Learning Rate 5e-5 1.5e-5
End Learning Rate 5e-6 1.5e-6
Warmup Steps 250 1000
Tokens 1B 5.4B
Appendix B Detailed results
In this section, we present detailed results on all considered datasets. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 14.
Tables 8 and 9 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 8 contains detailed results from training on C# and Python. Table 9 contains results on C averaged across three different subsets of C (for details about the construction of those subsets see Appendix F). Both tables show that SPLiCe outperforms the Baseline by a significant margin and is slightly better than or on par with Repo.
The main advantage of SPLiCe over the Repo approach is that it can be used also for non-structured data. Table 10 shows the detailed results of applying the SPLiCe on non-code data. Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
Table 11 shows that a simple artificial extension of example length via random concatenation of documents does not help.
Table 8: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in mentioned ways. For details about training please refer to Section 3.1.
Altered Train Data: C#
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
ArXiv 5.52 5.53 5.65 5.53
C 2.39 2.40 2.50 2.40
C++ 2.60 2.61 2.74 2.62
CUDA 2.46 2.48 2.65 2.49
C# 1.82 1.82 1.90 1.82
Common Lisp 3.41 3.44 3.72 3.42
Dart 2.14 2.16 2.31 2.16
Emacs Lisp 8.41 8.46 8.85 8.41
Erlang 2.80 2.80 2.95 2.81
Fortran 3.68 3.71 4.05 3.72
Go 1.94 1.95 2.06 1.96
Groovy 3.01 3.03 3.25 3.04
Haskell 3.33 3.35 3.58 3.35
Java 2.09 2.10 2.22 2.10
Pascal 3.61 3.62 3.80 3.60
Python 2.90 2.91 3.07 2.91
mean 3.26 3.27 3.46 3.27
Altered Train Data: Python
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
ArXiv 5.47 5.49 5.57 5.48
C 2.38 2.39 2.45 2.39
C++ 2.61 2.63 2.71 2.63
CUDA 2.41 2.43 2.56 2.44
C# 1.98 1.99 2.06 2.00
Common Lisp 3.23 3.28 3.47 3.27
Dart 2.15 2.17 2.27 2.17
Emacs Lisp 7.94 7.98 8.28 7.93
Erlang 2.70 2.71 2.82 2.71
Fortran 3.42 3.46 3.72 3.46
Go 1.95 1.96 2.03 1.96
Groovy 2.97 2.99 3.13 2.99
Haskell 3.25 3.28 3.46 3.27
Java 2.12 2.12 2.20 2.13
Pascal 3.57 3.58 3.73 3.58
Python 2.53 2.53 2.62 2.54
mean 3.17 3.19 3.32 3.18
Table 9: To check the statistical significance of our results we prepare three subsets of C (details in Appendix F and train the models on a 50/50 mixture of RedPajama data (organized in standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
Altered Train Data: C
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
Common Lisp 2.994 $\pm$ 0.072 3.042 $\pm$ 0.091 3.195 $\pm$ 0.078 3.043 $\pm$ 0.102
Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
Emacs Lisp 7.857 $\pm$ 0.017 7.851 $\pm$ 0.020 8.098 $\pm$ 0.027 7.840 $\pm$ 0.019
Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
Table 10: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods. For details about training please refer to Section 3.1. Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
Altered Train Data: StackExchange
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
ArXiv 5.07 5.09 5.14
C 2.68 2.69 2.70
C++ 3.02 3.04 3.06
CUDA 2.89 2.93 2.94
C# 2.27 2.28 2.29
Common Lisp 4.02 4.06 4.08
Dart 2.58 2.60 2.61
Emacs Lisp 9.55 9.67 9.69
Erlang 3.13 3.16 3.18
Fortran 4.28 4.34 4.38
Go 2.24 2.25 2.27
Groovy 3.62 3.66 3.68
Haskell 3.88 3.91 3.94
Java 2.43 2.45 2.45
Pascal 4.08 4.11 4.14
Python 3.32 3.35 3.36
mean 3.69 3.73 3.74
Altered Train Data: Wikipedia
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
ArXiv 5.64 5.65 5.73
C 2.65 2.67 2.71
C++ 2.98 3.01 3.07
CUDA 2.87 2.92 3.00
C# 2.22 2.24 2.29
Common Lisp 3.87 3.96 4.08
Dart 2.51 2.55 2.61
Emacs Lisp 9.38 9.45 9.63
Erlang 3.13 3.16 3.23
Fortran 4.23 4.32 4.49
Go 2.18 2.21 2.26
Groovy 3.49 3.55 3.67
Haskell 3.82 3.87 3.97
Java 2.39 2.41 2.46
Pascal 4.32 4.23 4.40
Python 3.26 3.30 3.37
mean 3.68 3.72 3.81
Table 11: Perplexity results comparing different ways of organizing the same data. All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C code is organized in one of three ways. For details about training please refer to Section 3.1.
Altered Train Data: C
Eval/Method SPLiCe BM25 Baseline Random
ArXiv 5.46 5.55 5.55
C 2.13 2.17 2.18
C++ 2.40 2.47 2.47
CUDA 2.22 2.33 2.33
C# 1.94 2.02 2.02
Common Lisp 2.99 3.20 3.18
Dart 2.14 2.27 2.27
Emacs Lisp 7.86 8.10 8.09
Erlang 2.67 2.77 2.77
Fortran 3.31 3.55 3.56
Go 1.91 2.00 2.00
Groovy 3.01 3.15 3.15
Haskell 3.20 3.37 3.37
Java 2.07 2.16 2.16
Pascal 3.49 3.62 3.64
Python 2.81 2.93 2.93
Mean 3.10 3.23 3.23
Appendix C Ablations
C.1 SPLiCe parameters
There are two important design choices related to SPLiCe. First, how many related documents are retrieved in each step (the parameter $k$ in Algorithm 1). Second, how the documents are ordered. Table 12 indicates that $k=1$ is the best choice, though the differences are rather small. We found that changing the order of documents in training examples hardly matters. We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling. We consider exploring these design choices to be an interesting future work direction.
Table 12: Ablation of SPLiCe hyper-parameters. For each ablation, we have trained the same $270M$ parameter model using different data organization methods. Top-$k$ corresponds to the number of descendants chosen in the $\texttt{RETRIEVE}(d,k)$ step of the Algorithm 1. Reverse and shuffle correspond to the final order of examples $C$ returned by the Algorithm 1 (reverse – the order of documents in $C$ is reversed, shuffle – examples are shuffled.
Code
Method Top-$k$ Order C++ Haskell Python CUDA All
SPLiCe BM25 top 1 standard 2.38 3.20 2.82 2.23 2.93
reverse 2.38 3.21 2.82 2.23 2.93
top 2 standard 2.39 3.23 2.84 2.24 2.95
reverse 2.39 3.24 2.84 2.24 2.95
top 3 standard 2.39 3.24 2.84 2.25 2.97
reverse 2.39 3.25 2.84 2.25 2.96
shuffle 2.40 3.24 2.84 2.25 2.96
We also evaluated the influence of BOS and EOS tokens on the performance of trained models. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively slightly decrease the number of separating tokens compared to the Baseline. However, in Appendix C.2 we included experiments showing that this has no impact on model performance.
C.2 Importance of separating tokens
We also evaluate the influence of BOS and EOS tokens on the performance. To be more precise, in all setups training examples are separated by BOS and EOS tokens. As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens. To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement). The results are shown in Table 13. We note that the difference between the Baseline and the random concatenation approach is small and the random concatenation approach does not result in significant perplexity gains.
Table 13: Perplexity evaluation of two methods of organizing the data. Baseline – document equals training example. Random – concatenate documents into examples of length bounded by $120$k characters. Training examples are then fed into the model and separated by BOS and EOS tokens. The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context. Experiments were performed using $270$M parameter model. We performed three runs on different subsets of C provide mean and standard deviation.
Training Method arXiv Code Code &
Data Python All arXiv
C Random 5.554 $\pm$ 0.004 2.931 $\pm$ 0.003 3.076 $\pm$ 0.005 3.231 $\pm$ 0.005
Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
Appendix D Key-Value retrieval task
Figure 4 shows how training on SPLiCe organized data improves the performance on the key-value retrieval task proposed in (Liu et al., 2023). This is a zero-shot task in which a model is prompted with a JSON dictionary and asked to retrieve a value corresponding to a specified key. The structure of the input is showcased below.
Extract the value corresponding to the specified key in the JSON object below.
JSON data:
{"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",
 "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",
 "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",
 "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",
 "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",
 "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",
 "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",
 "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}
 "4701aa05-c523-4b89-9700-64ab9c37c537":
Appendix E Perplexity Improvements
In Figure 5 we present perplexity improvements of SPLiCe BM25 over Baseline. Figure 6 shows evolution of SPLiCe model perplexity during the training. We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets. We average preplexity across arXiv, CUDA, Haskell and CommonCrawl datasets.
Figure 5: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps). We bucket tokens by their positions in the document, and calculate the average. Each dot is the difference of the averages of the SPLiCe and Baseline models. We observe that SPLiCe has smaller perplexity, and the improvements tend to be larger for tokens further in the document.
Figure 6: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens. See 5 for the difference with the baseline. As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence. Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
Appendix F Data preparation
F.1 Evaluation Data
We have taken a random subset of arXiv from Proof-pile. For StarCoder data, we have downloaded up to $64$GB of each of the mentioned language subsets and performed a random 85/15 split for languages that we train on.
When evaluating the model we skip documents that are shorter than $32$K tokens and truncate documents that are longer than that. Table 14 shows the number of tokens over which the perplexity was calculated.
Table 14: Number of evaluation tokens in each of the considered datasets.
Eval Dataset Number of Tokens
ArXiv 16M
C 16M
C# 16M
Common Lisp 16M
C++ 16M
CUDA 14M
Dart 16M
Emacs Lisp 15M
Erlang 16M
Fortran 16M
Go 16M
Groovy 4M
Haskell 16M
Java 16M
Pascal 16M
Python 16M
F.2 Train Data
The StackExchange data was taken from the Proof-pile. To prepare the code train data, we take the StarCoder train splits mentioned in Section F.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs. We have decided to divide the dataset in this way to not weaken the Repo method (a simple random split would scatter repository files across the data packs). We also split repos larger than $25$MB and filter out files that are longer than $30$k characters. The reason behind repo splitting is to avoid the situation where one repository occupies a significant portion of the data pack. We have noticed repos containing as many as $40$k files and files as long as $11$M characters. The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data. For C# and Python, only one pack is used for Baseline, Repo, and SPLiCe to organize the data. For C we have performed a run on three packs and provided results and standard deviation in Table 2. For the $3B$ model we run the methods on several packs and concatenate the results into a single dataset.
Appendix G TREC accuracy improvements
Figure 7 show details about accuracy improvements on TREC when considering different numbers of in-context examples.
(a) Examples: $190$, Context $4K$
(b) Examples: $380$, Context $8K$
(c) Examples: $780$, Context $16K$
(a) Examples: $190$, Context $4K$
(b) Examples: $380$, Context $8K$
(c) Examples: $780$, Context $16K$
(d) Examples: $1560$, Context $32K$
Figure 7: Histograms of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task. The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples. Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC. Note that the model trained with SPLiCe is almost always better than the Baseline.
Appendix H Faiss parameters
Our experiments with SPLiCe Cont utilize Faiss (Johnson et al., 2017a) for fast approximate inner-product search. To be more precise we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.

### v5

@@
 Abstract
-Recent advances in long-context Large Language Models (LCLMs) have generated significant interest, especially in applications such as querying scientific research papers.
-However, their potential is often limited by inadequate context utilization.
-We identify the absence of long-range semantic dependencies in typical training data as a primary hindrance.
-To address this, we delve into the benefits of frequently incorporating related documents into training inputs.
-Using the inherent directory structure of code data as a source of training examples, we demonstrate improvements in perplexity, even for tasks unrelated to coding.
-Building on these findings, but with a broader focus, we introduce Structured Packing for Long Context (SPLiCe).
-SPLiCe is an innovative method for creating training examples by using a retrieval method to collate the most mutually relevant documents into a single training context.
-Our results indicate that SPLiCe enhances model performance and can be used to train large models to utilize long contexts better.
-We validate our results by training a large $3$B model, showing both perplexity improvements and better long-context performance on downstream tasks.
+Recent developments in long-context large language models have attracted considerable attention.
+Yet, their real-world applications are often hindered by ineffective context information use.
+This work shows that structuring training data to increase semantic interdependence is an effective strategy for optimizing context utilization.
+To this end, we introduce Structured Packing for Long Context (SPLiCe), a method for creating training examples by using information retrieval methods to collate mutually relevant documents into a single training context.
+We empirically validate SPLiCe on large $3$B and $7$B models, showing perplexity improvements and better long-context utilization on downstream tasks.
+Remarkably, already relatively short fine-tuning with SPLiCe is enough to attain these benefits.
+Additionally, the comprehensive study of SPLiCe reveals intriguing transfer effects such as training on code data leading to perplexity improvements on text data.
+Machine Learning, Language Models, ICML
 1 Introduction
-Large language models (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023; Bai et al., 2023) have completely transformed the field of AI and natural language processing.
-Long-context language models (LCLMs) are becoming increasingly popular and have the potential to unlock new capabilities (Anthropic, 2023).
-However, recent studies (Liu et al., 2023; Tworkowski et al., 2023) have highlighted numerous issues with their context utilization.
-Even if they have the technical ability to process long context, some models fail in practice, as evidenced by their poor performance even on synthetic retrieval tasks (Li et al., 2023a).
-Moreover, LLMs are often distracted from retrieving relevant information when provided with multiple documents in the prompt (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023), and typically struggle to utilize information from the middle of their input (Liu et al., 2023).
-There might be multiple reasons behind these issues, requiring structured research to address them.
+Large language models (LLMs) (Brown et al., 2020; Chowdhery et al., 2022; Lewkowycz et al., 2022; OpenAI, 2023b; Bai et al., 2023) have transformed the field of AI.
+Recently, the field has observed the rise of Long-Context Language Models (LCLMs) that promise to unveil novel and powerful capabilities (Anthropic, 2023; OpenAI, 2023a).
+However, their ability to process long context is not always as effective as one hopes for.
+Indeed, several studies have highlighted an important limitation: when processing prompts composed of multiple documents, LCLMs frequently encounter difficulties in accurately extracting relevant information (Tworkowski et al., 2023; Liu et al., 2023; Shi et al., 2023a).
+Additionally, they typically find it challenging to utilize information from the middle of their inputs (Liu et al., 2023), even on simple synthetic retrieval tasks (Li et al., 2023a).
+Understanding these issues is vital for advancements in LCLM technologies and calls for systematic research.
 In this work, we take a step towards better context utilization in LCLMs.
-We focus on training data, keeping other components (e.g., the architecture and training objectives) unchanged.
-The fundamental question then is how to construct training examples such that using long context is beneficial for the model (e.g., it allows for obtaining lower perplexity or acquiring new downstream capabilities) while being non-trivial (like simple copying).
-We show that including multiple related documents in a training context can be beneficial - not only for evaluating on multiple documents, but also when considering language modeling perplexity on long, naturally occurring, single documents, which our evaluation is aimed for.
-First, we showcase this using the natural directory structure of the code data.
-Following that, we propose SPLiCe, Structured Packing for Long Context, that constructs training examples consisting of multiple similar documents selected using a retrieval method of choice (e.g., BM25).
-Figure 1: Training examples generated by the standard sampling procedure and proposed SPLiCe method.
-We empirically validate SPLiCe on a large-scale code generation dataset (Li et al., 2023b), demonstrating improved perplexity evaluated at larger contexts, as well as a higher performance on TREC (Li & Roth, 2002; Hovy et al., 2001), Qasper, and synthetic task from (Liu et al., 2023).
-For that, we use the dataset generated by our method to continue pretraining an OpenLLaMA 3B LLM using the methodology described in (Tworkowski et al., 2023), resulting in a long-context language model with improved capabilities.
+We focus on training data, keeping other components, such as the architecture and training objectives, unchanged.
+The broad question is how to organize training data to enhance long context capabilities?
+Such perspective has received some attention recently (Levine et al., 2022; Chan et al., 2022; Shi et al., 2023b), but the problem is far from being solved.
+The central finding of this work is that structuring training data to increase semantic interdependence is an effective strategy towards better long context utilization.
+We achieve this by introducing and evaluating Structured Packing for Long Context (SPLiCe), a method for creating training examples by using retrieval (e.g., BM25 or repository structure) to collate mutually relevant documents into a single training context.
+Figure 1: Training examples generated by the standard sampling procedure vs proposed by SPLiCe.
+Similar colors indicate related documents, which could be found using a retrieval method (e.g., BM25 or contriver) or metadata (e.g., repository structure for code).
+We empirically validate SPLiCe showing that fine-tuning of OpenLLaMA $3$Bv2 and $7$Bv2 (Geng & Liu, 2023) for only $2$B–$6$B tokens already brings perplexity reduction.
+This reduction translates to substantial improvements in handling long-context information in downstream tasks that require retrieval and in-context learning.
+These tasks include TREC (Li & Roth, 2002; Hovy et al., 2001), DBpedia (Lehmann et al., 2015), Qasper (Shaham et al., 2022; Dasigi et al., 2021), HotPotQA (Yang et al., 2018), and the lost-in-the-middle benchmark (Liu et al., 2023).
+We perform a comprehensive study of the design choices and properties of SPLiCe, showing that the acquired long-context capabilities transfer between modalities, such as code and text.
 Our contributions can be summarized as follows:
-  • We show that structuring training data is a viable way of improving the context utilization of LLMs. We validate this using the natural directory structure of code data.
-  • We introduce SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes retrieval (implemented using BM25 or Contriever (Izacard et al., 2022)) to find relevant documents without relying on directory structure.
-We show that using SPLiCe during training on code matches the model’s performance of the directory-structure method while being more generally applicable.
-  • We present a thorough analysis of the design choices of SPLiCe, the number of retrieved documents, and the order in which documents are merged into a training example.
-  • We validate SPLiCe by training a large-scale model, OpenLLaMA 3B v2 (Geng & Liu, 2023).
-We show that SPLiCe leads to improved perplexity on long-context evaluation, improved in-context learning ability and better retrieval performance.
+  • We comprehensively show that structuring training data is a viable way of improving the long context utilization of LLMs.
+To this end, we introduce SPLiCe, a method for creating training examples by using retrieval to collate mutually relevant documents into a single training context.
+  • We fine-tune OpenLLaMA $3$Bv2 and OpenLLaMA $7$Bv2 (Geng & Liu, 2023) using SPLiCe, showing that it improves perplexity on long context evaluation, in-context learning ability, and ability to retrieve relevant data from the context.
+  • We present a thorough analysis of the design choices behind SPLiCe, such as the number of retrieved documents and the order in which documents are merged into a training example.
 2 Method
-SPLiCe is a method of constructing the training examples that allows improvement of large-scale models’ long-context utilization and, consequently, their performance on a range of tasks.
-Through experimental study, we positively validate that such an approach leads to better long-context capabilities.
-Rationale and intuitions
-Capturing long-range dependencies is believed to enhance language modeling, e.g.
-as demonstrated in the study by (Borgeaud et al., 2022) which highlights the advantages of incorporating global dependencies in retrieval-augmented language models.
-However, in typical datasets, the occurrence of such dependencies is rare (de Vries, 2023) and diminishes as the distance between elements increases, leading to challenges in model learning.
-Our approach addresses this challenge by focusing on increasing dependency density, achieved by creating training examples that consist of multiple related documents.
-The rationale of this strategy is reinforced by the work of Levine et al.
-(2022), which shows that the trained model establishes stronger dependencies between text segments present in the same training example.
-In our work, we study the following methods:
-Baseline
-The standard approach, commonly used in LLM training pipelines (Brown et al., 2020), is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context (also known as example packing).
-Repository-level code data
-Certain datasets encompass meta-information regarding the data structure, which can be utilized to devise training examples.
-Take, for instance, a dataset consisting of GitHub repositories (e.g., (Li et al., 2023b)): rather than grouping documents randomly into a single context, we might group documents based on their subdirectory in a repository, thereby crafting a training example with potentially more cross-document dependencies compared to those generated randomly.
-This methodology, elaborated in (Wu et al., 2022), led to notable perplexity enhancements with augmented context length, suggesting that data orchestrated in such a manner could exhibit valuable long-context training signals.
-In our study, we strive to quantify the advantages of this methodology against the random packing baseline outlined above while maintaining the set of training documents to eliminate potential confounders.
-Figure 2: Visualization of SPLiCe training example generation.
-Structured Packing for Long Context (SPLiCe)
-The additional structure provided by meta-information is unavailable in most datasets.
-We propose to overcome this challenge by using a retrieval method to identify related documents.
-We test two choices of retrieval methods BM25 (Robertson & Zaragoza, 2009; Bassani, 2023) and Contriever-MSMARCO (Izacard et al., 2022)).
-With those tools at our disposal, we propose SPLiCe, a method that creates training examples by building a tree of documents.
-It starts from a randomly sampled document and continues in a breadth-first manner, expanding each node with top-$k$ similar documents from the corpus.
-The final sequence consists of a flattened tree according to a certain traversal strategy (e.g., level-order traversal).
-An intuitive visualization of SPLiCe is presented in Figures 1 and 2, while a detailed pseudocode is given in Algorithm 1.
-We note that the general formulation of SPLiCe is flexible.
-For example, the parameter $k$ allows for interpolation between two modes of retrieval.
-For $k=1$ SPLiCe creates a path of related examples simulating a long document.
-For larger values of $k$ our method generates examples that are similar to the ones used by retrieval augmented models.
-Note that the dataset must be reasonably curated for SPLiCe to work efficiently.
-For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
+SPLiCe is a method for constructing training examples that enhances the LLMs long-context utilization.
+This translates to improved performance on a range of tasks like in-context learning, question answering, in-context information retrieval, and long context language modeling.
+We make the source code available to ensure the reproducibility of our results.
+(For this anonymous submission we share the zip file with our code.)
 Algorithm 1 SPLiCe training example construction
-Input:
-   $D$: the document corpus,
-   $k$: the breadth hyper-parameter
-   $L$: Maximum length of the returned training example
+  Input:    $D$: document corpus   $k$: breadth hyper-parameter   $L$: maximum length of returned training example   RETRIEVE: retrieval method to use, e.g., BM25   ORDER: ordering method, e.g., id, or random shuffle
   Output: training example consisting of concatenated documents
-$d_{r}\sim D$triangleright Sample the root document
+  $d_{r}\sim D$Sample the root document}
   $D=D\setminus\{d_{r}\}$
-$C=d_{r}$
+  $C=[d_{r}]$
   $Q=\texttt{empty queue }$
@@
      $d=Q.\texttt{POP}()$
-      $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$triangleright Retrieve top $k$ most similar documents to $d$ using a                                                         selected method, e.g., BM25
+     $d_{1},\ldots,d_{k}=\texttt{RETRIEVE}(d,k)$
+     {Retrieve top-$k$ most similar documents to $d$ using a selected method, e.g., BM25}
      for each $d_{i}$ in $d_{1},\ldots,d_{k}$ do
-         if $d_{i}\in D$ then $\triangleright$ RETRIEVE uses a precomputed index, and may return                                                     documents that are already in $C$
-                  $C=C.\texttt{CONCAT}(d_{i})$triangleright Append $d_{i}$ to $C$
+        if $d_{i}\in D$ then
+           {RETRIEVE uses a precomputed index, and may return documents that are already in $C$}
+           $C=C.\texttt{APPEND}(d_{i})$Append $d_{i}$ to $C$}
            $Q.\texttt{PUSH}(d_{i})$
            $D=D\setminus\{d_{i}\}$
-return $\texttt{TRIM}(C,L)$
+        end if
+     end for
+  end while
+  return $\texttt{TRIM}(\texttt{ORDER}(C),L)$
+Rationale and intuitions
+Capturing long-range dependencies is believed to enhance language modeling and retrieval-augmentation (Borgeaud et al., 2022).
+However, it is an open question how to achieve such benefits in pre-training or fine-tuning.
+The primary difficulty comes from the fact that long-range dependencies are rare in training data (de Vries, 2023) and diminish with distance.
+It is thus unlikely that a model will learn to utilize long context without more explicit guidance.
+Recent studies indicate that structuring data, i.e., going beyond the i.i.d.
+paradigm, might be beneficial or even necessary to achieve good long-context performance.
+(Levine et al., 2022) develops a theory showing that the trained model establishes stronger dependencies between text segments in the same training example.
+(Chan et al., 2022) indicates that specific distributional properties of the training data have a significant impact on the model’s in-context performance.
+Finally, (Shi et al., 2023b) shows that training on structured data improves the performance of LLMs on long-context tasks.
+SPLiCe follows these intuitions.
+Namely, it constructs training examples by collating mutually relevant documents to increase the dependency density, thus allowing the model to learn to utilize long context.
+We provide additional insights into the distributional properties of SPLiCe created data at the end of this section and in Appendix K.
+Structured Packing for Long Context (SPLiCe)
+We propose SPLiCe, a method that creates training examples by building a tree of documents, see Algorithm 1 and Figure 1 for a general overview.
+SPLiCe starts by picking a random document from the dataset to create a single-node tree and continues in a breadth-first manner, each time appending top-$k$ similar documents from the corpus.
+The final sequence consists of the tree flattened according to a certain tree traversal strategy.
+The hyperparameter $k$ introduces flexibility and allows to interpolate between modes of retrieval.
+Indeed, for $k=1$ SPLiCe creates a path of related examples simulating a long document.
+For larger values of $k$, SPLiCe produces examples that are similar to the ones used by retrieval augmented models, e.g., (Borgeaud et al., 2022).
+Note that the dataset must be reasonably curated for SPLiCe to work efficiently.
+For example, a low duplication rate incentivizes the model to utilize long context beyond simple copying.
+Many possible retrieval methods can be used with SPLiCe (Retrieve function in Algorithm 1). In our experiments, we test the following:
+  • SPLiCeRepo: based on additional meta-information about the data, that is the repository structure of the code (Repo): we concatenate files using a depth-first search algorithm on the directory structure, that is files from the same directory are grouped together.
+A similar method has been pioneered by (Wu et al., 2022) and proposed in (Shi et al., 2023b) as an interesting future direction.
+  • SPLiCeBM25: based on BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), a standard retrieval method that uses a bag-of-words approach to rank documents based on their similarity to a query.
+  • SPLiCeCont: based on Contriever-MSMARCO (Cont) (Izacard et al., 2022), a recently proposed retrieval method that uses a transformer model to rank documents based on their similarity to a query.
+SPLiCe computational efficiency Given the dataset sizes used in training LLMs, the computational efficiency plays a non-trivial role.
+SPLiCeRepo is the fastest and easy to implement but requires additional directory structure, i.e., is not applicable to general web data.
+SPLiCeBM25 requires quadratic time index and needs dataset chunking to be applicable to large datasets, see Section 3.3.
+SPLiCeCont requires calculating embeddings for each document and retreval based on the cosine similarity.
+The latter step can be done efficiently using fast approximate inner-product search, e.g., Faiss (Johnson et al., 2017).
+Baseline The standard approach, commonly used in LLM training pipelines, is to randomly sample documents from the dataset and concatenate them to make training examples of a desired context.
+This is known as example packing (Brown et al., 2020).
+Burstiness SPLiCe conjecturally falls into the framework presented in (Chan et al., 2022).
+This paper shows that the distributional properties of the training data affect the in-context capabilities of transformer models.
+In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence .
+In Appendix K, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) compared to the baseline.
 3 Experiments with Medium-Scale Models
-In this section, we conduct a comprehensive examination of the impact of document packing into training examples on long-context performance.
-We demonstrate that leveraging the inherent repository structure of code data can yield notable perplexity enhancements.
-Additionally, we establish that our proposed method, SPLiCe, either matches or surpasses this approach, exhibiting broader applicability to non-code data.
-We stress that we do not aim to surpass the approach to concatenate repository files together, instead we aim to show that SPLiCe is more general and can be used for non-code data.
-We carry out an ablation study on SPLiCe to scrutinize its design elements and hyperparameters in Appendix C.1.
-For this segment, due to computational constraints, we engage with $270$M models to ensure a meticulous investigation.
-In Section 4, we affirm that these findings extend to large-scale models.
+In this section, we conduct a comprehensive examination of the impact of document packing on long-context performance, using medium-scale models (see Section 4 for large models results).
+We demonstrate that structuring data with SPLiCe is a generally applicable approach to improve long context performance.
+We study the performance in Section 3.2 and design choices of SPLiCe in Section 3.3.
 3.1 Experimental Setup
-To investigate the impact of input data organization on enabling language models to leverage extended context, we employ $270$M models following a specific protocol.
-Initially, we pre-train a $270$M model on $6.3$B tokens from the original RedPajama mixture (TogetherComputer, 2023), using a context size of $2$K.
-Subsequently, we continue pretraining this model on $1$B tokens with an extended context length of $32$K, employing the Focused Transformer (FoT) objective Tworkowski et al.
-(2023), on a balanced mixture of the original RedPajama data (TogetherComputer, 2023) and long-context data assembled using the specified method.
-This approach is driven by practical factors: training with a shorter context expedites the process, and as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023), the model training can be continued on longer documents from the original data source while maintaining the desired long-context perplexity scaling.
-During the long-context pretraining phase, only the latter half of the mixture is altered, retaining the original RedPajama portion intact.
-This approach is loosely inspired by (Ouyang et al., 2022) and (Rozière et al., 2023) and aims to prevent the model from overfitting to artificially created long documents.
-Only the effect of data organization into training examples is tested, keeping other factors like architecture intact.
-For a detailed description of the model architecture, refer to Appendix A, and for information regarding the training data, see Appendix F.
-In the evaluation, we employ a context length of $32$K and measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets.
-The selection of these evaluation datasets is motivated by their inclusion of documents that stand to benefit from long-context information.
-Specifically, functions are often re-utilized, and similar terms are employed across papers.
-The advantage of long context in these datasets has been demonstrated for both arXiv (Chen et al., 2023) and StarCoder (Li et al., 2023b).
+In this part, we employ $270$M models and the following protocol.
+Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023).
+Subsequently, we continue training using $1$B tokens with an extended context length of $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline.
+We employ the Focused Transformer (FoT) objective (Tworkowski et al., 2023) for context extension (unless stated otherwise).
+This approach is motivated by practical factors, viz.
+training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023).
+Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture.
+This aims to prevent the model from overfitting to artificially created long documents.
+In the evaluation, we measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K.
+The selection of these datasets is motivated by their inclusion of documents that can benefit from long-context information as demonstrated in (Chen et al., 2023) and (Li et al., 2023b).
+For example, functions are often re-utilized, and similar terms are employed across papers.
 We exclude documents with fewer than $32$K tokens and truncate those exceeding this length.
-For further details regarding the chosen datasets and the evaluation setup, refer to Appendix F.
-3.2 SPLiCe on code data
-In this section, we compare SPLiCe against the Baseline and repository-level code data (Repo) packing methods.
-We test two retrieval approaches for SPLiCe, the first one is BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), the second one is and Contriever-MSMARCO (Izacard et al., 2022)) and denote them as SPLiCe BM25 and SPLiCe Cont respectively.
-In SPLiCe BM25 retrieval takes into account the whole content of the document, whereas in SPLiCe Cont we use the first $512$ tokens and Faiss (Johnson et al., 2017b) for fast approximate inner-product search (for details see Appendix H).
-To make this comparison possible we use code data (required by Repo).
-We observe, see Table 1, that SPLiCe significantly outperforms the Baseline, and is either slightly better or on par with Repo.
-This is a promising result, as SPLiCe is more general and can be applied to non-code data (see Section 3.3).
-Table 1: We train a $270$M parameter model on a 50/50 mixture of the RedPajama data (organized in a standard way) and code data (organized in one of four ways: SPLiCe BM25, SPLiCe Cont, Repo, Baseline – standard).
-We evaluate the perplexity of the models on arXiv and various subsets of StarCoder.
-The boldface denotes the best result within the training mixture.
-Note that SPLiCe is generally on par or better than Repo and beats the Baseline by a large margin.
-For detailed results, see Appendix B.
-Training Method arXiv Code Code &
+We refer to Appendix A for a detailed description of the model architecture. For information regarding the training and evaluation data, see Appendix G.
+3.2 Experimental Results
+In this section, we compare SPLiCe against Baseline.
+We test two retrieval approaches for SPLiCe: BM25 (Robertson & Zaragoza, 2009; Bassani, 2023), and Contriever-MSMARCO (Izacard et al., 2022)) denoted as SPLiCeBM25 and SPLiCeCont respectively.
+In SPLiCeBM25 retrieval takes into account the whole content of the document, whereas in SPLiCeCont we use the first $512$ tokens and Faiss (Johnson et al., 2017) for fast approximate inner-product search (see Appendix H for details).
+By SPLiCeRepo, we denote the approach using the directory structure of code repositories.
+We also test our method with $64$K context length and report the results in Table 19 in Appendix J.
+In those experiments, we use $k=1$; see Section 3.3 for more details.
+The results are presented in Table 3, Table 3, and Table 3, from which we draw the following conclusions.
+Structuring data improves performance.
+We observe (see Table 3) that all variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset.
+This indicates that the structure of the data is important for long-context training.
+The positive effects transfer between domains.
+We observe (see Table 3) positive effects when training on code and evaluating on arXiv and when training on web data and evaluating on code.
+This indicates that the positive effects of SPLiCe transfer between domains.
+We conjecture that better data mixtures could bring further benefits.
+Retrieval method is of secondary importance.
+We observe (see Table 3) that the choice of retrieval method has a minor impact on the results.
+Typically, SPLiCeBM25 is slightly better than SPLiCeCont.
+On the code data, SPLiCeRepo is on par with SPLiCeBM25.
+We recall that SPLiCeRepo can be applied only to code data, whereas SPLiCeBM25 and SPLiCeCont are more general.
+In practice, we often observe that SPLiCeCont is faster than SPLiCeBM25 due to the use of Faiss.
+[…]
+Long Context Method arXiv Code Code &
 Data Haskell Python CUDA All arXiv
-C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
-SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
+C SPLiCeBM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
+SPLiCeCont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
+SPLiCeRepo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
 Baseline 5.55 3.37 2.93 2.33 3.07 3.23
-Repo 5.47 3.22 2.83 2.24 2.96 3.12
-C# SPLiCe BM25 5.52 3.33 2.90 2.46 3.11 3.26
-SPLiCe Cont 5.53 3.35 2.91 2.48 3.12 3.27
+C# SPLiCeBM25 5.52 (.13) 3.33 (.25) 2.90 (.17) 2.46 (.19) 3.11 (.20) 3.26 (.20)
+SPLiCeCont 5.53(.12) 3.35(.23) 2.91(.16) 2.48(.17) 3.12(.19) 3.27(.19)
+SPLiCeRepo 5.53 (.12) 3.35 (.23) 2.91 (.16) 2.49 (.16) 3.12 (.19) 3.27 (.19)
 Baseline 5.65 3.58 3.07 2.65 3.31 3.46
-Repo 5.53 3.35 2.91 2.49 3.12 3.27
-Python SPLiCe BM25 5.47 3.25 2.53 2.41 3.02 3.17
-SPLiCe Cont 5.49 3.28 2.53 2.43 3.03 3.19
+Python SPLiCeBM25 5.47 (.10) 3.25 (.21) 2.53 (.09) 2.41 (.15) 3.02 (.15) 3.17 (.15)
+SPLiCeCont 5.49 (.08) 3.28 (.18) 2.53 (.09) 2.43 (.13) 3.03 (.14) 3.19 (.13)
+SPLiCeRepo 5.48 (.09) 3.27 (.19) 2.54 (.08) 2.44 (.12) 3.03 (.14) 3.18 (.14)
 Baseline 5.57 3.46 2.62 2.56 3.17 3.32
-Repo 5.48 3.27 2.54 2.44 3.03 3.18
-To deepen the analysis, we prepared three subsets of the C code and ran the tested methods on each of them.
-In Table 2, we report the mean and standard deviation.
-We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
-For details about how the training sets were prepared please refer to Appendix F.
-Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. For the C code we use three different subsets and report the mean and standard deviation.
-Training Method arXiv Code Code &
+Wikipedia SPLiCeBM25 5.64 (.09) 3.82 (.15) 3.26 (.11) 2.87 (.13) 3.55 (.13) 3.68 (.13)
+SPLiCeCont 5.65 (.08) 3.87 (.10) 3.30 (.07) 2.92 (.08) 3.59 (.09) 3.72 (.09)
+Baseline 5.73 3.97 3.37 3.00 3.68 3.81
+StackExchange SPLiCeBM25 5.07 (.07) 3.88 (.06) 3.32 (.04) 2.89 (.05) 3.60 (.05) 3.69 (.05)
+SPLiCeCont 5.09 (.05) 3.91 (.03) 3.35 (.01) 2.93 (.01) 3.63 (.02) 3.73 (.01)
+Baseline 5.14 3.94 3.36 2.94 3.65 3.74
+Table 2: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C code. We perform three trainings using different training datesets and report the mean and standard deviation.
+Long Context Method arXiv Code Code &
 Data Python All arXiv
-C SPLiCe BM25 5.463 $\pm$ 0.002 2.810 $\pm$ 0.002 2.942 $\pm$ 0.005 3.100 $\pm$ 0.004
-SPLiCe Cont 5.477 $\pm$ 0.005 2.824 $\pm$0.001 2.957 $\pm$ 0.006 3.115 $\pm$ 0.006
-Baseline 5.550 $\pm$ 0.002 2.931 $\pm$ 0.008 3.073 $\pm$ 0.006 3.228 $\pm$ 0.005
-Repo 5.474 $\pm$ 0.007 2.827 $\pm$ 0.006 2.958 $\pm$ 0.009 3.115 $\pm$ 0.009
-3.3 SPLiCe on webtext data
-The major advantage of our method over Repo is that SPLiCe can be applied to any textual data.
-Table 3 shows that using our method the model outperforms the baseline when trained on a $/50$ 50 mixture of RedPajama (packed in the standard way) and Wikipedia/StackExchange (prepared using SPLiCe).
-Notably, we observe that the model trained on StackExchange data significantly outperforms the one trained on C in the arXiv evaluation.
-Despite the non-code nature of StackExchange we still note perplexity improvements on code.
-Unsurprisingly, the perplexity on code is not as good as when training directly on code data.
-We conjecture that better data mixtures could bring further benefits.
-Table 3: Perplexity results for training on a $/50$ 50 data mixture of RedPajama and one of {StackExchange,C} organized using mentioned methods.
-Note that when evaluating on arXiv the model trained with SPLiCe organized StackExchange outperforms the one trained on C.
-Training Method arXiv Code Code &
-data Haskell Python CUDA All arXiv
-Wikipedia SPLiCe BM25 5.64 3.82 3.26 2.87 3.55 3.68
-SPLiCe Cont 5.65 3.87 3.30 2.92 3.59 3.72
-Baseline 5.73 3.97 3.37 3.00 3.68 3.81
-StackExchange SPLiCe BM25 5.07 3.88 3.32 2.89 3.60 3.69
-SPLiCe Cont 5.09 3.91 3.35 2.93 3.63 3.73
-Baseline 5.14 3.94 3.36 2.94 3.65 3.74
-C SPLiCe BM25 5.46 3.20 2.81 2.22 2.94 3.10
-SPLiCe Cont 5.48 3.22 2.82 2.23 2.96 3.11
-Baseline 5.55 3.37 2.93 2.33 3.07 3.23
-3.4 SPLiCe with different context extension approaches
-Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension.
+C SPLiCeBM25 5.463 $\pm$ .002 2.810 $\pm$ .002 2.942 $\pm$ .005 3.100 $\pm$ .004
+SPLiCeCont 5.477 $\pm$ .005 2.824 $\pm$.001 2.957 $\pm$ .006 3.115 $\pm$ .006
+SPLiCeRepo 5.474 $\pm$ .007 2.827 $\pm$ .006 2.958 $\pm$ .009 3.115 $\pm$ .009
+Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
+[…]
+RoPe scale Training Method arXiv Code Code &
+method data Haskell Python CUDA All arXiv
+Naive C# SPLiCeBM25 6.25 (.08) 4.84 (.19) 3.55 (.11) 2.84 (.12) 3.72 (.13) 3.88 (.12)
+SPLiCeRepo 6.25 (.08) 4.87 (.16) 3.56 (.10) 2.85 (.11) 3.74 (.11) 3.89 (.11)
+Baseline 6.33 5.03 3.66 2.96 3.85 4.00
+CodeLlama C# SPLiCeBM25 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.53 (.05) 3.34 (.06) 3.49 (.06)
+SPLiCeRepo 5.74 (.02) 4.28 (.09) 3.22 (.05) 2.54 (.04) 3.35 (.05) 3.50 (.05)
+Baseline 5.76 4.37 3.27 2.58 3.40 3.55
+YaRN C# SPLiCeBM25 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.55 (.06) 3.37 (.07) 3.52 (.06)
+SPLiCeRepo 5.77 (.02) 4.32 (.10) 3.24 (.05) 2.56 (.05) 3.38 (.06) 3.53 (.05)
+Baseline 5.79 4.42 3.29 2.61 3.44 3.58
+SPLiCe works with various context extension results Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023) for context extension.
 FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers.
@@
 Table 3 shows the results.
-Table 4: Perplexity results for training on a $/50$ 50 data mixture of RedPajama, and C# organized using the mentioned methods.
-We check how much SPLiCe can help when fine-tuning for the larger context using methods different than FoT (Tworkowski et al., 2023).
-That is, we test the effectiveness of SPLiCe when the context is extended to $16K$ in all layers and parameters of positional embedding are either adjusted as in CodeLlama (Rozière et al., 2023), with YaRN (Peng et al., 2023), or left without changes.
-The results indicate that SPLiCe still brings perplexity improvements in all of these setups.
-For YaRN we have used the parameters from (Peng et al., 2023) and adjusted the scale factor appropriately.
-RoPe scale Training Method arXiv Code Code &
-method data Haskell Python CUDA All arXiv
-Naive C# SPLiCe BM25 6.25 4.84 3.55 2.84 3.72 3.88
-Baseline 6.33 5.03 3.66 2.96 3.85 4.00
-Repo 6.25 4.87 3.56 2.85 3.74 3.89
-CodeLlama C# SPLiCe BM25 5.74 4.28 3.22 2.53 3.34 3.49
-Baseline 5.76 4.37 3.27 2.58 3.40 3.55
-Repo 5.74 4.28 3.22 2.54 3.35 3.50
-YaRN C# SPLiCe BM25 5.77 4.32 3.24 2.55 3.37 3.52
-Baseline 5.79 4.42 3.29 2.61 3.44 3.58
-Repo 5.77 4.32 3.24 2.56 3.38 3.53
+Training is stable.
+We prepared three subsets of the C code dataset and ran the tested methods on each of them.
+In Table 3, we report the mean perplexity and its standard deviation.
+We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
+3.3 Design Choices Analysis
+Choice of k As already mentioned in the previous section, the choice of the retrieval method (BM25 or Contriver) plays a minor role.
+In Appendix C.1, we study the role of $k$.
+We first observe that its choice has a minor impact on the results, with $k=1$ being the best choice.
+We leave for further work a more detailed analysis of this hyperparameter.
+Document ordering We discovered that the ordering of the documents collected by the retrieval procedure might be a subtle issue.
+We consider three choices of ORDER function in Algorithm 1: identity (do not permute the documents), reverse (reverse the order of documents) and a random shuffle.
+In Appendix C.1, we present results for $270$M models showing negligible differences.
+However, we observed differences for large models, elaborated in Section 4.2.4, in which random shuffling is the best choice for the CodeLlama context extension method.
+Corpus chunking Performing retrieval from the whole corpus might be cumbersome due to its scale.
+However, we find that chunking the corpus into smaller parts still brings improvements.
+In particular, for the experiments with English Wikipedia and BM25 presented in Table 3, we used a manageable size of $2$GB.
 4 Large-Scale Models
 In this section, we show that SPLiCe can improve the long context performance of large-scale language models.
-To this end, we use $3$B parameter models and tasks that test in-context learning, question answering, and information retrieval capabilities.
-We provide perplexity results in Appendix E
+To this end, we use $3$B and $7$B parameter models and tasks that test in-context learning (Section 4.2.1), question answering (Section 4.2.2), and in-context information retrieval capabilities (Section 4.2.3).
+We provide perplexity results in Appendix E.
+We also emphasize that those improvements occur during a relatively short, compared to pre-training, fine-tuning.
+To be more precise, $3$B models were tuned on $5.4$B tokens, whereas $7$B models were tuned on $2$B tokens.
+All of the models were pre-trained on $1$T tokens.
 4.1 Experimental Setup
-We continue training OpenLLaMA 3B v2 (Geng & Liu, 2023; Geng, 2023) on the dataset defined in Section 3.1 - that is, a 50/50 mixture of RedPajama prepared in a standard way and C prepared using a method of choice.
-Specifically, we train two models with $32K$ context length for the same number of tokens, 5.4B, which is $21$k training steps, on two datasets: one generated using SPLiCe, and the other one using the Baseline, see Section 2 for definitions.
-We use a batch size of 256K tokens per step, and learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023).
-We evaluate the perplexity on a mixture of long documents sourced from CommonCrawl, arXiv, and StarCoder (Li et al., 2023b).
-Otherwise, we follow the protocol described in Section 3.1.
-4.1.1 In context learning
-To evaluate the improvements of the model’s in-context learning ability we use TREC (Li & Roth, 2002; Hovy et al., 2001) question classification task.
-We provide the SPLiCe trained model with $190$ to $1560$ in-context examples and compare the prediction accuracy against the Baseline model.
-We sample the few-shot examples using 50 different random seeds and show the histogram of accuracy improvements in Figure 3 and additional results in Table 4 and Figure 7.
-Figure 3: Histogram of accuracy improvement of SPLiCe BM25 over Baseline on TREC question classification task.
-The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples.
-Each set of in-context examples consists of $1560$ elements (roughly $32$k tokens) randomly sampled (without replacement) from the training subset of TREC.
-4.2 Standardized Comparison Over Long Language Sequences
-We show that SPLiCe can improve model performance on Qasper (Dasigi et al., 2021; Shaham et al., 2022).
-Qasper tests the model’s ability to answer questions regarding research papers provided as a part of the input.
-The evaluated models were trained on a mixture of RedPajama prepared in a standard way and C prepared either in a standard way (Baseline) or using SPLiCe.
-Note that despite SPLiCe using only code data here, we still see improvements in score on non-code, see Table 5.
-Table 5: Zero-shot performance on the validation subset of Qasper (Dasigi et al., 2021) and few-shot performance on TREC.
-For Qasper we use the implementation from Language Model Evaluation Harness (Gao et al., 2021).
-In the Harness implementation, yes/no questions are evaluated separately from open questions.
-Note that despite using SPLiCe for code data only we still have improvements in non-code tasks.
-For TREC we average results across $50$ sets of in-context examples.
-Note that for TREC SPLiCe model is better for almost all choices of in-context examples (see Figures 3 and 7).
-Task Context length 3B Baseline 3B SPLiCe
-TREC 32K 73.9 $\pm$3.9 79.3 $\pm$ 2.9
-16K 68.9 $\pm$5.9 76.9 $\pm$ 3.1
-8K 66.5 $\pm$6.2 75.8 $\pm$ 3.5
-4K 65.1 $\pm$6.1 72.8 $\pm$ 4.9
-Qasper F1 32K 23.1 23.9
-4K 18.6 18.7
-4.3 Key Retrieval Performance
-Recent works indicated numerous issues related to long-context utilization.
-In particular, (Liu et al., 2023, Figure 7) proposed a benchmark retrieval task to measure the ability of language models to retrieve information from a long context.
+For $3$B experiments, we continue training OpenLLaMA $3$B v2 (Geng & Liu, 2023; Geng, 2023) on a 50/50 mixture of RedPajama prepared in the standard way and C prepared using SPLiCeBM25.
+For $7$B ones, we continue the training of OpenLLaMA $7$B v2 on 75/25/25 mixture of RedPajama (50) prepared in the standard way, StackExchange (25) and C (25) prepared using SPLiCeBM25.
+StackExchange is part of RedPajama (TogetherComputer, 2023), and C data come from StarCoder (Li et al., 2023b).
+We train with $32$K context length using $5.4$B tokens for $3$B models and $2$B for $7$B models (recall that OpenLLaMA models have the $2$K context length).
+We use a batch size of $256$K tokens per step and the learning rate of ${1.5}\mathrm{e}{-5}$ with linear warmup and cosine decay, following (Geng & Liu, 2023).
+Regarding the context extensions method we use FoT (Tworkowski et al., 2023), and the CodeLlama (CL) (Rozière et al., 2023) approach. More explicitly, we test six models:
+[…]
+where Baseline denotes the standard data preparation method. Hyperparameter details are located in Appendix A.
+4.2 Experimental Results
+We present perplexity measurements in Appendix E.
+In particular, following (Anthropic, 2023), we examine perplexity improvements in the function of the token position in the document.
+Our findings reveal that SPLiCe models demonstrate superior perplexity across all distances.
+Notably, the improvement in perplexity is more pronounced for tokens positioned later in the document, indicating enhanced language modeling capabilities.
+For the remainder of this section, we focus on testing the in-context learning, question answering and information retrieval abilities on our models using downstream tasks.
+4.2.1 In-Context Learning
+To evaluate the improvements of in-context learning abilities, we use TREC (Li & Roth, 2002; Hovy et al., 2001) and DBpedia (Lehmann et al., 2015), which are text classification tasks.
+For TREC, to test the long context utilization, we vary the number of in-context examples in $\{190,380,780,1560\}$ (which correspond roughly to $\{4,8,16,32\}$K context length).
+We sample these examples multiple times and measure classification accuracy on test set.
+In Table 4, we observe that SPLiCe improves the average performance across different context lengths and model sizes.
+We follow a very similar protocol for DBpedia, see Table 5, confirming that SPLiCe improves the context utilization.
+In Appendix I, we study the distribution of improvements with respect to the choice of the in-context examples, showing the domination of SPLiCe.
+In Table 4 and Table 5, we report mean improvement $\Delta$ along with $95\%$ bootstrap confidence intervals. See Appendix I for related histograms.
+Table 4: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001).
+For TREC, we average across $50$ sets of in-context examples for $3$B models and $10$ sets for $7$B models.
+We use $\Delta$ [confidence interval] to denote mean improvement and its $95\%$ bootstrap confidence intervals (see Appendix I).
+TREC
+Model
+Context
+(#examples)
+Baseline SPLiCe $\Delta$ [conf interv]
+[…]
+$16$K (780) 68.9 76.9 8.0 [6.9,9.3]{6.9,9.3]}6.9 , 9.3 ]
+$8$K (380) 66.5 75.8 9.3 [8.1,10.8]{8.1,10.8]}8.1 , 10.8 ]
+$4$K (190) 65.1 72.8 7.7 [6.5,9.0]{6.5,9.0]}6.5 , 9.0 ]
+[…]
+$16$K (780) 74.0 79.0 5.0 [3.4,6.0]{3.4,6.0]}3.4 , 6.0 ]
+$8$K (380) 72.1 76.2 4.1 [2.9,5.7]{2.9,5.7]}2.9 , 5.7 ]
+$4$K (190) 66.9 69.7 2.8 [1.5,4.2]{1.5,4.2]}1.5 , 4.2 ]
+[…]
+$16$K (780) 81.4 82.5 1.1 [0.2,1.6]{0.2,1.6]}0.2 , 1.6 ]
+$8$K (380) 76.6 76.5 -0.1 [1.0,1.1]{-1.0,1.1]}- 1.0 , 1.1 ]
+$4$K (190) 68.1 69.1 1.0 [0.4,2.9]{-0.4,2.9]}- 0.4 , 2.9 ]
+Table 5: Average performance on DBpedia (Lehmann et al., 2015).
+We average results across $40$ sets of in-context examples, with $\Delta$ [confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence interval.
+Due to the size of the evaluation dataset for each set of in-context examples, we subsample a subset of $500$ elements of the evaluation set.
+DBpedia
+Model
+Context
+(#examples)
+Baseline SPLiCe $\Delta$ [conf interv]
+[…]
+$16$K (190) 79.1 82.0 2.9 [2.5,3.4]{2.5,3.4]}2.5 , 3.4 ]
+[…]
+$16$K (190) 83.6 85.6 2.0 [1.5,2.5]{1.5,2.5]}1.5 , 2.5 ]
+[…]
+$16$K (190) 96.2 96.4 0.2 [0.0,0.3]{0.0,0.3]}0.0 , 0.3 ]
+Table 6: Average performance on HotPotQA (Yang et al., 2018).
+We average results across $7$ sets of in-context examples, with with $\Delta$[confidence interval] denoting mean improvement and its $95\%$ bootstrap confidence intervals.
+HotPotQA
+Model
+Context
+(#examples)
+Baseline SPLiCe $\Delta$ [conf interv]
+[…]
+[…]
+[…]
+4.2.2 Question Answering
+We also show that SPLiCe models have improved question answering capabilities.
+To this end, we use Qasper (Dasigi et al., 2021; Shaham et al., 2022) and HotPotQA (Yang et al., 2018).
+Qasper contains questions regarding research papers provided as a part of the input context.
+In Table 7, we observe that SPLiCe improves the two-shot performance.
+Similarly, we observe improvements on HotPotQA, see Table 6.
+We stress that in both cases, the results measure in-context capabilities as neither of the datasets was used during training.
+Table 7: Two-shot F1 performance on the validation subset of Qasper (Dasigi et al., 2021).
+For Qasper, we use the implementation from Language Model Evaluation Harness (Gao et al., 2021).
+Note that in the $3$B model case, despite using SPLiCe for code data only, we still have improvements in non-code tasks.
+QASPER
+Model
+Context length
+(#examples)
+Baseline SPLiCe
+[…]
+[…]
+[…]
+4.2.3 Key Retrieval Performance
+The key retrieval task introduced in (Liu et al., 2023) has become a routine diagnostic tool to study context capabilities.
 Specifically, the model is presented with a list of key-value pairs, which are $128$-bit UUIDs, and is tasked to retrieve the value for a given key.
-Even though very simple, the task proved to be challenging for many open-source language models.
-In Figure 4 we present the key retrieval performance of the models trained with SPLiCe and Baseline.
-The dictionary contains $75$ pairs, which is about 6k tokens.
-The performance depends heavily on the key position (which is also the case for other models, as indicated in (Liu et al., 2023)).
-The keys near the beginning are much harder to retrieve, and SPLiCe improves the accuracy more than two times.
-Figure 4: Key-value retrieval performance on a dictionary of $75$ key-value pairs ($\approx$ 6k tokens).
-The model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions.
-Both models achieve $100$% accuracy near the end because the desired key is in the local context (below $2048$ tokens).
+Even though very simple, the task proved to be challenging for many open-source language models, see Appendix D for the task details.
+In Figure 2, we present the key retrieval performance of $7$B models trained with CodeLlama (Rozière et al., 2023) context extension method.
+We note that despite relatively short tuning, SPLiCe significantly helps on hard-to-retrieve positions.
+We provide the results for FoT models in Appendix D.
+Figure 2: Key-value retrieval performance on a dictionary of $300$ key-value pairs ($\approx$ $24$K tokens).
+The $7$B CL model trained with SPLiCe achieves much higher accuracy on hard-to-retrieve positions.
 The details about this task can be found in Appendix D.
+Each position was evaluated using $500$ examples.
+4.2.4 Shuffling
+In our experiments with large $7$B CodeLlama models, we have noted slight improvements in handling long inputs when shuffling documents inside the context, that is, setting Order in the Algorithm 1 to be a random shuffle.
+We provide a comparison on TREC in Table 17.
 5 Related Work
-Some recent studies have focused on enhancing training examples to improve the performance of language models.
+There is an increasing number of works aiming to study the role of data in LLM training in detail.
 For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance.
-Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include ”intrinsic tasks”.
+Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks.
 They showed that this approach substantially boosts in-context learning.
 Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022).
+A recent work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities.
+Similarly, (Han et al., 2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
 Our methodology diverges from these works in several key ways.
@@
 This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code.
+The gradient-based method presented in (Han et al., 2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
+Concurrently to our research, (Shi et al., 2023b) showed a similar method for preparing the training data.
+However, in our work, we focus on tuning the models for long context, up to $32$K tokens, whereas (Shi et al., 2023b) focuses on training the models from scratch, with context length up to $8$K tokens.
+In addition to that, we provide ablations regarding context extension methods.
 6 Limitations and Future Work
 We show that structuring the training data is a viable way of improving the model’s performance.
-The presented method can viewed as a general framework for organizing the documents into training examples (i.e.
-to achieve the repository-level grouping, it suffices to replace the $\texttt{RETRIEVE}(d,k)$ method with one that returns files from the closest catalog in the repository).
+The presented method can viewed as a general framework for organizing the documents into training examples.
 This opens multiple further research avenues.
-In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Contriever-MSMARCO).
+Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo).
 However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations.
-We leave the study of how the choice of retriever (in particular, how the ratio of related and unrelated documents) affects performance for future work.
-Note that as BM25 is a text-statistic based retriever, it is not guaranteed that the documents returned by it are semantically related.
+We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work.
+Note that as BM25 is a syntactic retriever, it is not guaranteed that the documents returned by it are semantically related.
 Moreover, we have not tuned Contriever-MSMARCO, and it was provided only with a $512$ token long prefix of a document.
-Another, avenue for future work is to study the granularity of the pieces from which the training examples are constructed.
+In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval).
+However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
+Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed.
 In this work, we focus on the document-level granularity.
 However, it is possible to construct training examples from smaller pieces, such as paragraphs or sentences.
-Our method requires further studies to understand its scaling properties beyond the range presented in the paper.
-In particular, we leave scaling the context length to hundreds of thousands of tokens and experiments with models having more than $3$B parameters for future work.
-Our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023).
-This methodology is similar to the approach used in (Wu et al., 2022) and allows extension of the model context with relatively small (compared to other methods) computational cost.
-Briefly speaking, it boils down to choosing a subset of attention layers for the context extension.
-We have additionally tested three popular context extension methods on a medium-scale (Naive, YaRN, and CodeLlama).
-We leave studies of SPLiCe with other context extension methods as future work.
-One of the approaches to training long-context language models is to use long conversational data (Li et al., 2023a).
-This approach is complementary to our method.
-SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, and answers, etc.) to further increase the number of long-context examples.
+Scaling Our method requires further studies to understand its scaling properties beyond the range presented in the paper.
+Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023).
+Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe.
+However, we leave more detailed studies of the design choices of SPLiCe with other context extension methods to future work.
+Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a).
+This is complementary to our method.
+SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, StackExchange questions, answers, etc.) to further increase the number of long-context examples.
 We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
-In order to prepare the training data SPLiCe uses each document exactly once (we mask out each used document for further retrieval).
-However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
-Using highly correlated samples has the potential to result in training instability.
-However, during our experiments with StarCoder, Wikipedia, and StackExchange data we have noted no performance degradation.
-We leave the study of how SPLiCe integrates with different types of data for the future.
-In particular, in our studies used datasets were reasonably deduplicated.
+Data curation Using highly correlated samples has the potential to result in training instability.
+However, we noted no performance degradation during our experiments.
+We leave the study of how SPLiCe integrates with different data types for the future.
 7 Conclusions
 In this work, we present SPLiCe, a novel method of constructing training examples for long-context language models, which utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner.
-We show that SPLiCe improves the perplexity of on language modeling task and improves the performance on downstream tasks.
-We further show that SPLiCe can be used to train large-scale models, resulting in a model with improved long-context utilization.
-Besides its direct usefulness, we believe that our work indicates multiple interesting research directions on how to improve the performance of long-context language models with structured data.
-8 Reproducibility
-To ensure the reproducibility of our results we provide the details about how the data was prepared in Appendix F, the pseudocode of our method in Algorithm 1 and details about the trained models in Appendix A.
-We also attach the source code in the supplementary material.
+We show that SPLiCe improves the perplexity of the language modeling task and performance on downstream tasks.
+We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning.
+We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
+8 Impact Statement
+This paper presents work whose goal is to improve the long context utilization ability of language models through structuring training data.
+It bears standard risks associated with advancing the field of LLMs (such as misuse and generation of malicious/misleading content).
+For a detailed overview, we refer to (Bender et al., 2021; Weidinger et al., 2021).
+Acknowledgments
+We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC).
+We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265.
+Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
 Appendix A Architecture
@@
 Additionally, we use Rotary Position Embedding (Su et al., 2021).
-For context extension, we use Focused Transformer (Tworkowski et al., 2023).
+For context extension, we use Focused Transformer (FoT)(Tworkowski et al., 2023), CodeLlama (Rozière et al., 2023) (CL) and YaRN (Peng et al., 2023).
 Table 9 presents the details about both standard and long-context pretraining.
 Table 8: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
-Parameter/Model Size 270M 3B
-Vocab Size 32000 32000
-Embedding Size 1024 3200
-Num Attention Layers 12 26
-Num Attention Heads 8 32
-Head Size 128 100
-MLP Hidden Size 4096 8640
-FoT Context Extension Layers [6, 9] [6, 12, 18]
-Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B parameter OpenLLaMAv2 model (Geng, 2023).
-Stage Parameter/Model Size 270M 3B
-Pretraining Context 2K 2K
-Tokens 6.3B 1T
-Long Context Pretraining Context 32K 32K
-Batch Size 128 128
-Start Learning Rate 5e-5 1.5e-5
-End Learning Rate 5e-6 1.5e-6
-Warmup Steps 250 1000
-Tokens 1B 5.4B
-Appendix B Detailed results
-In this section, we present detailed results on all considered datasets.
-The training methodology is described in Section 3.1.
-Details about the number of tokens used for evaluation can be found in Table 14.
-Tables 8 and 9 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method.
+Parameter/Model Size 270M 3B 7B
+Vocab Size 32000 32000 32000
+Embedding Size 1024 3200 4096
+Num Attention Layers 12 26 32
+Num Attention Heads 8 32 32
+Head Size 128 100 128
+MLP Hidden Size 4096 8640 11008
+FoT Context Extension Layers [6, 9] [6, 12, 18] [8, 16, 24]
+Table 9: Training details.
+We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B parameter OpenLLaMAv2 model (Geng, 2023).
+Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al., 2023) and no-FoT to other methods (Naive, YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023)).
+Stage Parameter/Model Size 270M 3B 7B
+Pretraining Context 2K 2K 2K
+Tokens 6.3B 1T 1T
+Long Context Pretraining Context 32KFoT, 16Kno-FoT 32K 32K
+Batch Size 128FoT, 16no-FoT 128 128FoT, 32no-FoT
+Start Learning Rate 5e-5 1.5e-5 1.5e-5
+End Learning Rate 5e-6 1.5e-6 1.5e-6
+Warmup Steps 250 1000 1000FoT, 200no-FoT
+Tokens 1B 5.4B 2B
+Appendix B Detailed Results For Medium-Size Models
+In this section, we extend results presented in Section 3. The training methodology is described in Section 3.1. Details about the number of tokens used for evaluation can be found in Table 18.
+Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method.
 Table 10 contains detailed results from training on C# and Python.
-Table 9 contains results on C averaged across three different subsets of C (for details about the construction of those subsets see Appendix F).
-Both tables show that SPLiCe outperforms the Baseline by a significant margin and is slightly better than or on par with Repo.
-The main advantage of SPLiCe over the Repo approach is that it can be used also for non-structured data.
+Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G).
+Both tables show that SPLiCe outperforms the Baseline by a significant margin.
+The main advantage of SPLiCeBM25/SPLiCeCont over the SPLiCeRepo approach is that it can be used also for non-structured data.
 Table 12 shows the detailed results of applying the SPLiCe on non-code data.
 Note that training on non-code data allows us to improve the model perplexity on the arXiv dataset in comparison to the model trained on code.
+Table 13 considers models trained with YaRN (Peng et al., 2023), CodeLlama (Rozière et al., 2023) and Naive (no adjustment to RoPE) context extension methods.
 Table 14 shows that a simple artificial extension of example length via random concatenation of documents does not help.
 Table 10: Perplexity results comparing different ways of organizing the same data.
-All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in mentioned ways.
+All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and code organized in the mentioned ways.
 For details about training please refer to Section 3.1.
 Altered Train Data: C#
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
+Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
 ArXiv 5.52 5.53 5.65 5.53
@@
 Altered Train Data: Python
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
+Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
 ArXiv 5.47 5.49 5.57 5.48
@@
 Mean 3.17 3.19 3.32 3.18
-Table 9: To check the statistical significance of our results we prepare three subsets of C (details in Appendix F and train the models on a 50/50 mixture of RedPajama data (organized in standard way) and C data organized using one of the methods.
+Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods.
 Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
 Altered Train Data: C
-Eval/Method SPLiCe BM25 SPLiCe Cont Baseline Repo
+Eval/Method SPLiCeBM25 SPLiCeCont Baseline SPLiCeRepo
 ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
@@
 All runs started from the same $270M$ model with $2048$ context and were trained for $32$K context on a 50/50 mixture of RedPajama (organized in a standard way) and other data organized using one of the methods.
-For details about training please refer to Section 3.1.
+For details about training, please refer to Section 3.1.
 Note that the model trained with SPLiCe on StackExchange outperforms the one trained on code on arXiv evaluation, showing the benefits of SPLiCe’s applicability to non-code data.
@@
 Mean 3.68 3.72 3.81
+Table 13: Perplexity results comparing different ways of organizing the same data for non-FoT models.
+All runs started from the same $270M$ model with $2048$ context and were trained for $16$K context on a 50/50 mixture of RedPajama (organized in a standard way) and C# code is organized in one of three ways.
+Altered Train Data: C#
+Context $16$K: CodeLlama
+Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
+ArXiv 5.74 5.76 5.74
+C 2.66 2.70 2.66
+C++ 2.79 2.83 2.79
+CUDA 2.53 2.58 2.54
+C# 1.91 1.93 1.91
+Common Lisp 3.78 3.85 3.79
+Dart 2.28 2.33 2.28
+Emacs Lisp 8.29 8.41 8.30
+Erlang 3.57 3.64 3.58
+Fortran 3.93 4.01 3.95
+Go 1.99 2.03 2.00
+Groovy 2.95 3.01 2.96
+Haskell 4.28 4.37 4.28
+Java 2.31 2.35 2.31
+Pascal 3.67 3.72 3.67
+Python 3.22 3.27 3.22
+Mean 3.49 3.55 3.50
+Context $16$K: YaRN
+Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
+ArXiv 5.77 5.79 5.77
+C 2.68 2.72 2.68
+C++ 2.81 2.85 2.81
+CUDA 2.55 2.61 2.56
+C# 1.92 1.94 1.92
+Common Lisp 3.83 3.92 3.84
+Dart 2.30 2.36 2.30
+Emacs Lisp 8.37 8.52 8.38
+Erlang 3.60 3.67 3.61
+Fortran 3.97 4.06 3.99
+Go 2.00 2.04 2.01
+Groovy 2.98 3.03 2.98
+Haskell 4.32 4.42 4.32
+Java 2.33 2.37 2.33
+Pascal 3.69 3.75 3.69
+Python 3.24 3.29 3.24
+Mean 3.52 3.58 3.53
+Context $16$K: Naive
+Eval/Method SPLiCeBM25 Baseline SPLiCeRepo
+ArXiv 6.25 6.33 6.25
+C 2.88 2.96 2.89
+C++ 3.04 3.13 3.05
+CUDA 2.84 2.96 2.85
+C# 2.04 2.08 2.04
+Common Lisp 4.40 4.56 4.39
+Dart 2.50 2.60 2.51
+Emacs Lisp 9.25 9.46 9.25
+Erlang 3.98 4.10 4.00
+Fortran 4.56 4.79 4.59
+Go 2.14 2.21 2.16
+Groovy 3.27 3.39 3.28
+Haskell 4.84 5.03 4.87
+Java 2.52 2.60 2.53
+Pascal 4.05 4.20 4.10
+Python 3.55 3.66 3.56
+Mean 3.88 4.00 3.89
 Table 14: Perplexity results comparing different ways of organizing the same data.
@@
 Table 15 indicates that $k=1$ is the best choice, though the differences are rather small.
-We found that changing the order of documents in training examples hardly matters.
+We found that changing the order of documents in training examples hardly matters for $270$M models.
 We use ’standard’, as ordered by Algorithm 1, the reversed order, and random shuffling.
-We consider exploring these design choices to be an interesting future work direction.
 Table 15: Ablation of SPLiCe hyper-parameters.
@@
 We also evaluate the influence of BOS and EOS tokens on the performance.
-To be more precise, in all setups training examples are separated by BOS and EOS tokens.
-As both Repo and SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens.
-To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement).
+To be more precise, in all setups, training examples are separated by BOS and EOS tokens.
+As SPLiCe methods concatenate documents to create training examples, they effectively increase the average example length and decrease the number of separating tokens.
+To check whether those methods do not simply benefit from the reduction in the number of BOS and EOS tokens, we have trained a model on data prepared similarly as in SPLiCe, but instead of most matching documents $\texttt{RETRIEVE}(d,k)$ returned random documents from the dataset (sampling without replacement).
 The results are shown in Table 16.
-We note that the difference between the Baseline and the random concatenation approach is small and the random concatenation approach does not result in significant perplexity gains.
+We note that the difference between the Baseline and the random concatenation approach is small, and the random concatenation approach does not result in significant perplexity gains.
 Table 16: Perplexity evaluation of two methods of organizing the data.
@@
 The difference is negligible, which suggests that the extension of example length in Random does not help the model to utilize extended context.
-Experiments were performed using $270$M parameter model.
-We performed three runs on different subsets of C provide mean and standard deviation.
-Training Method arXiv Code Code &
+Experiments were performed using the $270$M parameter model.
+We performed three runs on different subsets of C to provide mean and standard deviation.
+Long Context Method arXiv Code Code &
 Data Python All arXiv
@@
 Extract the value corresponding to the specified key in the JSON object below.
-JSON data:
-{"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",
- "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",
- "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",
- "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",
- "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",
- "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",
- "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",
- "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}
- "4701aa05-c523-4b89-9700-64ab9c37c537":
+JSON data: {"4ef217b7-6bc0-48c6-af35-2765f1e730f3": "068192b7-16b1-40e0-8495-61c63f979d50",  "cd6b8bdc-bc6c-4490-acb4-bc187a2dccba": "7364a26e-289f-4968-93d3-b273e882bdee",  "7d057372-4ab8-4811-8110-658c3f19fff4": "3ad075c5-b567-4201-85a7-cb31a0c91540",  "c62e192d-45e6-4646-bb88-1529c73256c9": "f0411644-1f6d-42a6-8af8-f06da66efc77",  "06134e93-e158-490e-a66c-8e3b98e12735": "50a26a36-d832-450c-8d6e-a4cc3d0ec0ab",  "3286f978-4270-4b54-8bfa-540d7e0772e6": "075cc716-1836-4f90-9be3-53e3d4ec6585",  "4701aa05-c523-4b89-9700-64ab9c37c537": "49d86354-74c4-4256-9b3a-35e6e2b80d00",  "c8895805-e574-4f13-9fe5-89da1d8c4748": "cc91af7f-8509-4bdc-bad7-2646af68e6d2"}  "4701aa05-c523-4b89-9700-64ab9c37c537":
+We noted that FoT-trained models struggle with this task.
+This is probably due to the fact that they extend context only in a couple of layers, and the key-value retrieval task requires looking up and extracting a long sequence of letters and digits.
+Because of that, we evaluate FoT models with shorter dictionaries consisting of $75$ key-value pairs (around $6$K tokens) and show the results in Figures 3, and 3.
+For comparison, we also evaluate the $7$B CL model with this context length and show the results in Figure 3.
+Figure 3: Performance on a smaller version of key-value retrieval task from (Liu et al., 2023).
+We note that FoT models (a), (b) generally struggle to retrieve tokens that are only visible to a subset of layers with extended context.
+For comparison, we show the results with a model that has extended context in all layers (c) using CodeLlama (Rozière et al., 2023) method of context extension.
+Each position was evaluated using $500$ examples.
 Appendix E Perplexity Improvements
-In Figure 5 we present perplexity improvements of SPLiCe BM25 over Baseline.
-Figure 6 shows evolution of SPLiCe model perplexity during the training.
+In Figure 4 we present perplexity improvements of $3$B FoT SPLiCeBM25 over Baseline.
+Figure 5 shows the evolution of SPLiCe model perplexity during the training.
 We follow (Anthropic, 2023) and bucket perplexity by token positions in buckets of length $2^{i}$ up to $32768$, and then average within the buckets.
-We average preplexity across arXiv, CUDA, Haskell and CommonCrawl datasets.
+We average perplexity across arXiv, CUDA, Haskell, and CommonCrawl datasets.
 Figure 4: Perplexity improvement with SPLiCe against the Baseline of the final models (after $21$k training steps).
-We bucket tokens by their positions in the document, and calculate the average.
+We bucket tokens by their positions in the document and calculate the average.
 Each dot is the difference of the averages of the SPLiCe and Baseline models.
@@
 Figure 5: Evolution of the perplexity with SPLiCe, as the model is trained on more tokens.
-See 5 for the difference with the baseline.
+See 4 for the difference with the baseline.
 As expected, SPLiCe significantly improves perplexity for tokens whose positions are very distant in the sequence.
 Perplexity for more distant tokens improves more significantly compared to tokens in the beginning, early in the training.
-Appendix F Data preparation
-F.1 Evaluation Data
+Appendix F Shuffling
+Table 17: Average classification performance on TREC (Li & Roth, 2002; Hovy et al., 2001).
+We compare $7$B CL model trained on SPLiCe prepared data with different approaches to ordering the examples (different function Order in 1).
+SPLiCe-shuf denotes the model trained on data that shuffled the documents randomly in context, for SPLiCe-no-shuf Order was an identity function.
+We use $\pm$ to denote the standard deviation.
+We decided to stick with the model trained with random shuffling as it has slightly better long-context performance and lower standard deviation.
+TREC
+Model
+Context length
+(# of examples)
+SPLiCe-no-shuf SPLiCe-shuf
+$7$B CL $32$K (1560) 76.0 $\pm$2.5 76.6 $\pm$1.9
+$8$K (380) 77.1 $\pm$2.4 76.5 $\pm$1.8
+Appendix G Data Preparation
+G.1 Evaluation Data
 We have taken a random subset of arXiv from Proof-pile.
 For StarCoder data, we have downloaded up to $64$GB of each of the mentioned language subsets and performed a random 85/15 split for languages that we train on.
-When evaluating the model we skip documents that are shorter than $32$K tokens and truncate documents that are longer than that.
+When evaluating the perplexity of the model, we skip documents that are shorter than the model context and truncate documents that are longer than that.
 Table 18 shows the number of tokens over which the perplexity was calculated.
-Table 14: Number of evaluation tokens in each of the considered datasets.
-Eval Dataset Number of Tokens
-ArXiv 16M
-C 16M
-C# 16M
-Common Lisp 16M
-C++ 16M
-CUDA 14M
-Dart 16M
-Emacs Lisp 15M
-Erlang 16M
-Fortran 16M
-Go 16M
-Groovy 4M
-Haskell 16M
-Java 16M
-Pascal 16M
-Python 16M
-F.2 Train Data
+Table 18: Number of evaluation tokens in each of the considered datasets. For each context length $c$, we consider only documents having not less than $c$ tokens and extract the $c$ tokens prefix.
+Eval/Method 16K 32K 64K
+ArXiv 16M 16M 16M
+C 16M 16M 16M
+C++ 16M 16M 16M
+CUDA 16M 14M 8M
+C# 16M 16M 16M
+Common Lisp 16M 16M 16M
+Dart 16M 16M 16M
+Emacs Lisp 16M 15M 9M
+Erlang 16M 16M 12M
+Fortran 16M 16M 16M
+Go 16M 16M 16M
+Groovy 11M 4M 2M
+Haskell 16M 16M 15M
+Java 16M 16M 16M
+Pascal 16M 16M 16M
+Python 16M 16M 16M
+G.2 Train Data
 The StackExchange data was taken from the Proof-pile.
-To prepare the code train data, we take the StarCoder train splits mentioned in Section F.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs.
-We have decided to divide the dataset in this way to not weaken the Repo method (a simple random split would scatter repository files across the data packs).
+To prepare the code train data, we take the StarCoder train splits mentioned in Section G.1, shuffle them, group the documents by the repository (documents from the same repository occur one after another), and split them into smaller packs.
 We also split repos larger than $25$MB and filter out files that are longer than $30$k characters.
@@
 The character filtering is consistent with our method as we aim to improve the performance in a scenario that lacks high-quality long-context data.
-For C# and Python, only one pack is used for Baseline, Repo, and SPLiCe to organize the data.
-For C we have performed a run on three packs and provided results and standard deviation in Table 2.
-For the $3B$ model we run the methods on several packs and concatenate the results into a single dataset.
-Appendix G TREC accuracy improvements
-Figure 7 show details about accuracy improvements on TREC when considering different numbers of in-context examples.
-(a) Examples: $190$, Context $4K$
-(b) Examples: $380$, Context $8K$
-(c) Examples: $780$, Context $16K$
+For C# and Python, only one pack is used to organize the data.
+For C, we have performed a run on three packs and provided results and standard deviation in Table 3.
+For large models, we run the methods on several packs and concatenate the results into a single dataset.
+For natural language datasets, we extract a random subset of documents.
+Appendix H Faiss Parameters
+Our experiments with SPLiCeCont utilize Faiss (Johnson et al., 2019) for fast approximate inner-product search.
+To be more precise, we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
+Appendix I Detailed Accuracy Improvements
+The performance of in-context learning depends much on the choice of the in-context examples. To study this in more detail we study the following random variable
+$\Delta(c)=\text{ACC}_{\textsc{{SPLiCe}}}(c)-\text{ACC}_{\textsc{Baseline}}(c),$
+where $\text{ACC}_{\textsc{{SPLiCe}}}(c),\text{ACC}_{\textsc{Baseline}}(c)$ are the accuracies of the model trained with SPLiCe and Baseline respectively on the random choice of in-context examples $c$.
+Below we report the histograms of $\delta(c)$.
+In Table 4 and Table 5 we report mean $\Delta$ and $95\%$ confidence intervals as $\Delta$ [confidence interval].
+Figure 6 shows details about accuracy improvements on TREC when considering different numbers of in-context examples.
 (a) Examples: $190$, Context $4$K
@@
 Figure 6: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on TREC question classification task.
-The results are obtained by comparing the accuracy on the test set of TREC of the 3B model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples.
+The results are obtained by comparing the accuracy on the test set of TREC of the $3$B FoT model trained with SPLiCe to the model trained with default data preparation method (Baseline) across $50$ sets of in-context examples.
 Each set of in-context examples consists of elements randomly sampled (without replacement) from the training subset of TREC.
 Note that the model trained with SPLiCe is almost always better than the Baseline.
-Appendix H Faiss parameters
-Our experiments with SPLiCe Cont utilize Faiss (Johnson et al., 2017a) for fast approximate inner-product search.
-To be more precise we use the ”IVF8192,Flat” index that we train on 262144 examples coming from the dataset.
+(a) Examples: $190$, Context $16$K, Model: $3$B FoT
+(b) Examples: $380$, Context $32$K, Model: $3$B FoT
+(c) Examples: $190$, Context $16$K, Model: $7$B FoT
+(d) Examples: $380$, Context $32$K, Model: $7$B FoT
+(e) Examples: $190$, Context $16$K, Model: $7$B CL
+(f) Examples: $380$, Context $32$K, Model: $7$B CL
+Figure 7: Histograms of accuracy improvement of SPLiCeBM25 over Baseline on DBPedia.
+We sample $40$ sets of in-context examples and, for each set of in-context examples, evaluate on a random $500$ element subset of the DBPedia test set.
+Appendix J Longer Context
+In Table 19, we present the perplexity results for $270$M parameter models trained as described in Section 3 but with 2x larger context extension – context extended from $2$K to $64$K, instead of from $2$K to $32$K.
+[…]
+Long Context Method arXiv Code Code &
+Data Haskell Python CUDA All arXiv
+C# SPLiCeBM25 4.86 (.15) 2.60 (.19) 2.66 (.16) 2.32 (.19) 2.75 (.19) 2.88 (.19)
+SPLiCeCont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
+SPLiCeRepo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
+Baseline 5.01 2.79 2.82 2.51 2.94 3.07
+Appendix K Data Distribution Property
+Chan et al.
+(2022); Han et al.
+(2023) shows that the ”burstiness” property of training data benefits a model’s generalisation.
+Specifically, ”burstiness” presents a flatter frequency distribution, with a relatively higher mass on the rare, long-tail tokens appearing in a sequence, which results in a lower Zipf’s coefficient of token frequency.
+We also observe a burstiness property of data in SPLiCe, as shown in Table 20.
+Table 20: Zipf’s coefficient of token frequency on Baseline and SPLiCe. A lower Zipf’s coefficient represents a more significant burstiness property.
+Training Data Context Length Method Zipf’s Coefficient
+C 2K SPLiCeBM25 ${1.667}_{(0.139)}$
+Baseline ${1.717}_{(0.152)}$
+StackEx 2K SPLiCeBM25 ${1.836}_{(0.088)}$
+Baseline ${1.865}_{(0.074)}$
+C 32K SPLiCeBM25 ${1.512}_{(0.055)}$
+Baseline ${1.593}_{(0.025)}$
+StackEx 32K SPLiCeBM25 ${1.643}_{(0.026)}$
+Baseline ${1.664}_{(0.013)}$